### v1

# A determinantal formula for orthosymplectic Schur functions

Abstract.
We prove a new determinantal formula for the characters of irreducible representations of orthosymplectic Lie superalgebras analogous to the formula developed by Moens and Jeugt (J. Algebraic Combin., 2003) for general linear Lie superalgebras. Our proof uses the Jacobi–Trudi type formulas for orthosymplectic characters. As a consequence, we show that the odd symplectic characters introduced by Proctor (Invent. Math., 1988) are the same as the orthosymplectic characters with some specialized indeterminates. We also give a generalization of an odd symplectic character identity due to Brent, Krattenthaler and Warnaar (J. Combin. Theory Ser. A, 2016).
Key words and phrases:
orthosymplectic Schur functions, hook Schur polynomials, determinantal formula
2010 Mathematics Subject Classification:
05E05, 05E10
1. Introduction
Hook Schur polynomials (or supersymmetric Schur polynomials) and orthosymplectic Schur polynomials are the characters of irreducible representations of general linear and orthosymplectic Lie superalgebras respectively. These are supersymmetric functions in two sets of commuting variables indexed by integer partitions. It is well known that hook Schur polynomials form the basis for the ring of supersymmetric functions. For background on the representation theory of Lie superalgebras, see [CW12].
Hook Schur polynomials were introduced by Berele and Regev [BR87] combinatorially using hybrid tableaux involving a column-strict part and a row-strict part. Stembridge [Ste85] gave a different characterization for these polynomials and showed that they form a basis for the ring of supersymmetric functions. Moens and Jeugt [MvdJ03] proved the following determinantal formula for these polynomials using the representation theory of general linear superalgebra $\mathfrak{gl}(m/n)$.
Theorem 1.1 ([MvdJ03, Equation 1.17]).
Let $\lambda$ be a partition such that $\lambda_{n+1}\leq m$ and $k=\min\{i|\lambda_{i}+n+1-i\leq m\}.$ Then the hook Schur polynomial $\operatorname{hs}_{\lambda}(X/Y)$ is given by
$\operatorname{hs}_{\lambda}(X/Y)=\frac{(-1)^{mn-n+k-1}}{D}\det\left(\begin{array}[]{c|c}\left(\frac{1}{x_{i}+y_{j}}\right)_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq m\end{subarray}}&\left(x_{i}^{\lambda_{j}+n-m-j}\right)_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq k-1\end{subarray}}\\ \\ \hline\cr\\ \left(y_{j}^{\lambda^{\prime}_{i}+m-n-i}\right)_{\begin{subarray}{c}1\leq i\leq m-n+k-1\\ 1\leq j\leq m\end{subarray}}&{\text{\huge 0}}\end{array}\right),$
where
$D=\frac{\prod_{1\leq i<j\leq n}(x_{i}-x_{j})\prod_{1\leq i<j\leq m}(y_{i}-y_{j})}{\prod_{i=1}^{n}\prod_{j=1}^{m}(x_{i}+y_{j})}.$
The main result of this article is to prove the following determinantal formula for orthosymplectic Schur polynomials analogous to the formula in 1.1 for hook Schur polynomials.
Theorem 1.2.
Let $\lambda$ be a partition of length at most $n$ and $k=\min\{j|\lambda_{j}+n+1-j\leq m\}.$ Then the orthosymplectic Schur function $\operatorname{spo}_{\lambda}(X/Y)$ is given by
(1.1) $\operatorname{spo}_{\lambda}(X/Y)=\frac{(-1)^{mn-n+k-1}}{D}\det\left(\begin{array}[]{c|c}R&X_{\lambda}\\[-8.5359pt] \\ \hline\cr\\[-8.5359pt] Y_{\lambda}&0\end{array}\right)$
where
$R=\left(\frac{x_{i}}{(x_{i}+y_{j})\prod_{q=1}^{m}(\bar{x}_{i}+y_{q})}-\frac{\bar{x}_{i}}{(\bar{x}_{i}+y_{j})\prod_{q=1}^{m}(x_{i}+y_{q})}\right)_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq m\end{subarray}},$
$X_{\lambda}=\left(\frac{x_{i}^{\lambda_{j}+n-m-j+1}}{\prod_{q=1}^{m}(\bar{x}_{i}+y_{q})}-\frac{\bar{x}_{i}^{\lambda_{j}+n-m-j+1}}{\prod_{q=1}^{m}(x_{i}+y_{q})}\right)_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq k-1\end{subarray}},\qquad Y_{\lambda}=\left(y_{j}^{\lambda^{\prime}_{i}+m-n-i}\right)_{\begin{subarray}{c}1\leq i\leq m-n+k-1\\ 1\leq j\leq m\end{subarray}}$
and
$D=\frac{\prod_{i=1}^{n}(x_{i}-\bar{x}_{i})\prod_{1\leq i<j\leq n}(x_{i}+\bar{x}_{i}-x_{j}-\bar{x}_{j})\prod_{1\leq i<j\leq m}(y_{i}-y_{j})}{\prod_{i=1}^{n}\prod_{j=1}^{m}(x_{i}+y_{j})(\bar{x}_{i}+y_{j})}.$
Example 1.3.
Let $n=1$, $m=2$ and consider the partition $\lambda=(2)$. Then $k=2$ and by (1.1), we have
$\operatorname{spo}_{\lambda}(x_{1}/y_{1},y_{2})=\frac{(x_{1}+y_{1})(x_{1}+y_{2})(\bar{x}_{1}+y_{1})(\bar{x}_{1}+y_{2})}{(x_{1}-\bar{x}_{1})(y_{1}-y_{2})}\\ \times\left(\frac{x_{1}}{(\bar{x}_{1}+y_{1})(\bar{x}_{1}+y_{2})}-\frac{\bar{x}_{1}}{(x_{1}+y_{1})(x_{1}+y_{2})}\right)\times\det\left(\begin{array}[]{cc}y_{1}&y_{2}\\ 1&1\end{array}\right).$
It turns out to be equal to
$x_{1}^{2}+\bar{x}_{1}^{2}+1+y_{1}(x_{1}+\bar{x}_{1})+y_{2}(x_{1}+\bar{x}_{1})+y_{1}y_{2}.$
Corollary 1.4.
Let $\lambda$ be a partition of length at most $n$. Then
$\operatorname{spo}_{\lambda}(x_{1},x_{2},\dots,x_{n}/y)=\frac{1}{D}\det\left(x_{i}^{\lambda_{j}+n-j+1}-\bar{x}_{i}^{\lambda_{j}+n-j+1}+y\left(x_{i}^{\lambda_{j}+n-j}-\bar{x}_{i}^{\lambda_{j}+n-j}\right)\right),$
where $D=\prod_{i=1}^{n}(x_{i}-\bar{x}_{i})\prod_{1\leq i<j\leq n}(x_{i}+\bar{x}_{i}-x_{j}-\bar{x}_{j})$.
We prove 1.2 in Section 3 after giving all required definitions in Section 2. Our proof uses the Jacobi–Trudi type formula for orthosymplectic Schur polynomials. The Jacobi–Trudi type formula was given by Balantekin and Bars [BB81]. It was proved using a different approach by Benkart, Shader and Ram [BSR98] and recently, Stokke and Visentin [SV16] constructed the formula using lattice paths. Orthosymplectic Cauchy identities and Pieri rule for orthosymplectic Schur polynomials were proved in [PPS22] and  [Sto18] respectively.
We also show that the odd symplectic characters introduced by Proctor [Pro88] are the same as the orthosymplectic Schur polynomials with some specialized indeterminates (see 2.11). Lastly, we give a generalization of an odd symplectic character identity due to Brent, Krattenthaler and Warnaar, which was found in the study of discrete Mehta-type integrals (see  [BKW16, BKW15, Oka20]) in Section 4.
2. Preliminaries
Recall that a partition $\lambda=(\lambda_{1},\lambda_{2},\dots,\lambda_{\ell})$ is a sequence of weakly decreasing nonnegative integers. Any entry of the partition is called a part. The length of a partition $\lambda$, which is the number of positive parts, is denoted by $\ell(\lambda)$. A partition $\lambda$ can be represented pictorially as a Young diagram, whose $i$’th row contains left-justified boxes. We will use the so-called English notation where the first row is on top. The conjugate of a partition $\lambda$, denoted $\lambda^{\prime}$, is the partition whose Young diagram is obtained by transposing the Young diagram of $\lambda$.
Throughout, we fix $n$ and $m$ to be positive integers. Let $X=(x_{1},x_{2},\dots,x_{n})$ and $Y=(y_{1},\dots,y_{m})$ be tuples of commuting indeterminates. Define $\bar{x}=1/x$ for an indeterminate $x$ and write $\overline{X}=(\bar{x}_{1},\dots,\bar{x}_{n})$. We use the shorthand notation $[a]$ to denote the set $\{1,2,\dots,a\}$ for $a\in\mathbb{Z}_{+}$.
We consider the ring $\mathbb{Z}[x_{1},\dots,x_{n}]$ of polynomials in $n$ commuting variables $x_{1},\dots,x_{n}$ with integer coefficients. A polynomial in this ring is symmetric if it is invariant under the action of permuting the variables. The elementary symmetric function $e_{\lambda}(X)$ and the complete symmetric function $h_{\lambda}(X)$ indexed by a partition $\lambda=(\lambda_{1},\lambda_{2},\dots,\lambda_{n})$ are defined as:
(2.1) $e_{\lambda}(X)=\prod_{i=1}^{n}e_{\lambda_{i}}(X),$
where
$e_{r}(X)=\begin{cases}\displaystyle\sum_{1\leq i_{1}<i_{2}<\dots<i_{r}\leq n}x_{i_{1}}x_{i_{1}}\dots x_{i_{r}}&r\geq 1,\\ \qquad\qquad\qquad\qquad 1&r=0,\end{cases}$
and
(2.2) $h_{\lambda}(X)=\prod_{i=1}^{n}h_{\lambda_{i}}(X),$
where
$h_{r}(X)=\begin{cases}\displaystyle\sum_{1\leq i_{1}\leq i_{2}\leq\dots\leq i_{r}\leq n}x_{i_{1}}x_{i_{1}}\dots x_{i_{r}}&r\geq 1,\\ \qquad\qquad\qquad\qquad 1&r=0.\end{cases}$
Also, define $e_{r}(X)=h_{r}(X)=0$ for $r<0$.
For a partition $\lambda=(\lambda_{1},\ldots,\lambda_{n})$ of length at most $n$, the Schur polynomial is given by the following determinantal formula:
(2.3) $s_{\lambda}(X)=\frac{\displaystyle\det_{1\leq i,j\leq n}\Bigl{(}x_{i}^{\lambda_{j}+n-j}\Bigr{)}}{\displaystyle\det_{1\leq i,j\leq n}\Bigl{(}x_{i}^{n-j}\Bigr{)}}.$
The expression in (2.3) is essentially the Weyl character formula. The following Jacobi–Trudi identity expresses the Schur polynomial $s_{\lambda}(X)$ in terms of complete symmetric functions,
(2.4) $s_{\lambda}(X)=\det\left(h_{\lambda_{i}-i+j}(X)\right)_{1\leq i,j\leq r},$
where $r$ is any integer such that $r\geq\ell(\lambda)$. Recall that there is a combinatorial formula for $s_{\lambda}(X)$ as a sum of monomials summed over all column-strict Young tableaux of shape $\lambda$. A column-strict Young tableau of shape $\lambda$ is a filling of $\lambda$ with entries from the set $[n]$ such that
  • the entries increase weakly along rows
  • the entries increase strictly along columns.
For partitions $\lambda$ and $\mu$ such that $\mu\subset\lambda\,(\mu_{i}\leq\lambda_{i}$, for all $i\geq 1)$, the skew shape $\lambda/\mu$ is the set theoretic difference $\lambda/\mu$ and the skew Schur function $s_{\lambda/\mu}(X)$ is given by
(2.5) $s_{\lambda/\mu}(X)=\sum_{T}\operatorname{wt}(T),\qquad\operatorname{wt}(T)\coloneqq\prod_{i\geq 1}x_{i}^{c_{i}(T)},$
where the sum is taken over all column-strict tableaux of shape $\lambda/\mu$ and $c_{i}(T)$, $i\in[n]$ is the number of occurrences of $i$ in $T$. Otherwise $s_{\lambda/\mu}(X)=0$. The following Jacobi–Trudi formula expresses $s_{\lambda/\mu}(X)$ in terms of complete symmetric functions.
(2.6) $s_{\lambda/\mu}(X)=\det\left(h_{\lambda_{i}-\mu_{j}-i+j}(X)\right)_{1\leq i,j\leq r},$
where $r$ is any integer such that $r\geq\ell(\lambda)$.
We now give a supersymmetric analogue of the symmetric functions. We consider the ring $\mathbb{Z}[x_{1},\dots,x_{n},y_{1},\dots,y_{m}]$ of polynomials in $(n+m)$ commuting variables $x_{1},\dots,x_{n},$ $y_{1},\dots,y_{m}$ with integer coefficients. A polynomial $f(X,Y)$ in this ring is doubly symmetric if it is symmetric in both $(x_{1},\dots,x_{n})$ and $(y_{1},\dots,y_{m})$. Moreover, if substituting $x_{n}=t$ and $y_{m}=-t$ results in an expression independent of $t$, then we call $f(X,Y)$ a supersymmetric function. The elementary supersymmetric function and the complete supersymmetric function indexed by $\lambda=(\lambda_{1},\lambda_{2},\dots,\lambda_{n})$ are:
(2.7) $E_{\lambda}(X/Y)=\prod_{i=1}^{n}E_{\lambda_{i}}(X/Y),$
where
$E_{r}(X/Y)=\begin{cases}\displaystyle\sum_{j=0}^{r}e_{j}(X)h_{r-j}(Y)&r\geq 1,\\ \qquad\qquad 1&r=0,\end{cases}$
and
(2.8) $H_{\lambda}(X/Y)=\prod_{i=1}^{n}H_{\lambda_{i}}(X/Y),$
where
$H_{r}(X/Y)=\begin{cases}\displaystyle\sum_{j=0}^{r}h_{j}(X)e_{r-j}(Y)&r\geq 1,\\ \qquad\qquad 1&r=0.\end{cases}$
It is convenient to define $E_{r}(X/Y)$ and $H_{r}(X/Y)$ to be zero for $r<0$.
Definition 2.1.
A supertableau $T$ of shape $\lambda$ with entries
$1<2<\cdots<n<1^{\prime}<2^{\prime}<\cdots<m^{\prime}$
is a filling of the shape with these entries satisfying the following conditions:
  • the entries increase weakly along rows and columns
  • the unprimed entries strictly increase along columns
  • the primed entries strictly increase along rows
The weight of such a supertableau is given by
$\operatorname{wt}(T)\coloneqq\prod_{i=1}^{n}x_{i}^{n_{i}(T)}\prod_{j=1^{\prime}}^{m^{\prime}}y_{j}^{n_{j}(T)},$
where $n_{\alpha}(T)$, for $\alpha$ being any of the entries given in 2.1, is the number of occurrences of $\alpha$ in $T$. Denote the set of supertableaux of skew shape $\lambda$ with entries given in 2.1 by $\operatorname{SSYT}_{n/m}(\lambda)$. For an integer partition $\lambda$, the hook Schur (or supersymmetric Schur) polynomial, denoted $\operatorname{hs}_{\lambda}(X/Y)$ is given by
(2.9) $\operatorname{hs}_{\lambda/\mu}(X/Y)\coloneqq\sum_{T\,\in\,\operatorname{SSYT}_{n/m}(\lambda)}\operatorname{wt}(T).$
Example 2.2.
Let $n=2$ and $m=1$ and consider the partition $(2,1)$. Then we have the following supertableaux in $\operatorname{SSYT}_{2/1}((2,1))$:
[…]
Therefore,
$\operatorname{hs}_{(2,1)}(x_{1},x_{2}/y_{1})=x_{1}^{2}x_{2}+x_{1}x_{2}^{2}+x_{1}^{2}y_{1}+2x_{1}x_{2}y_{1}+x_{2}^{2}y_{1}+x_{1}y_{1}^{2}+x_{2}y_{1}^{2}.$
The Jacobi–Trudi formula expresses $\operatorname{hs}_{\lambda}(X/Y)$ in terms of the complete supersymmetric functions [PT92].
(2.10) $\operatorname{hs}_{\lambda}(X/Y)=\det\left(H_{\lambda_{i}-i+j}(X/Y)\right)_{1\leq i,j\leq n}.$
Note that the hook Schur polynomial $\operatorname{hs}_{\lambda}(X/Y)$ is nonzero if and only if $\lambda_{n+1}\leq m$. Now we consider orthosymplectic Schur functions after defining symplectic Schur functions.
Definition 2.3.
A symplectic tableau $T$ of shape $\lambda$ with entries
$1<\overline{1}<2<\overline{2}<\cdots<n<\overline{n}$
is a filling of the shape with these entries satisfying the following conditions:
  • the entries increase weakly along rows and strictly along columns
[…]
The weight of such a tableau is given by
$\operatorname{wt}(T)=\prod_{i=1}^{n}x_{i}^{n_{\overline{i}}(T)-n_{i}(T)},$
where $n_{i}(T)$ and $n_{\overline{i}}(T)$ is the number of occurrences of $i$ and $\overline{i}$ in $T$ respectively. Denote the set of symplectic tableaux of shape $\lambda$ filled with entries given in 2.3 by $\operatorname{SPT}_{n}(\lambda)$.
Example 2.4.
The following is a symplectic tableau $T$ of shape $\lambda=(3,2,2)$ with $\operatorname{wt}(T)=x_{2}$.
{ytableau}$\ytableau 1&\overline{1}2\\ \overline{2}\overline{2}\\ 4\overline{4}$
The symplectic Schur function is defined as
$\mathrm{sp}_{\lambda}(X)\coloneqq\sum_{T\in\operatorname{SPT}_{n}(\lambda)}\operatorname{wt}(T),$
where the sum runs over the set of tableaux in $\operatorname{SPT}_{n}(\lambda)$. The symplectic Schur function can be expressed using the Weyl character formula in the following way. Let $\lambda$ be a partition of length at most $n$. Then
(2.11) $\mathrm{sp}_{\lambda}(X)=\frac{\displaystyle\det_{1\leq i,j\leq n}\Bigl{(}x_{i}^{\lambda_{j}+n-j+1}-\bar{x}_{i}^{\lambda_{j}+n-j+1}\Bigr{)}}{\displaystyle\det_{1\leq i,j\leq n}\Bigl{(}x_{i}^{n-j+1}-\bar{x}_{i}^{n-j+1}\Bigr{)}},$
and the denominator here is
(2.12) $\det_{1\leq i,j\leq n}\Bigl{(}x_{i}^{n-j+1}-\bar{x}_{i}^{n-j+1}\Bigr{)}=\prod_{i=1}^{n}(x_{i}-\bar{x}_{i})\,\prod_{1\leq i<j\leq n}(x_{i}+\bar{x}_{i}-x_{j}-\bar{x}_{j}).$
Definition 2.5.
Let $B_{0}=\{1,\overline{1},2,\overline{2},\dots,n,\overline{n}\}$, $B_{1}=\{1^{\prime},\dots,m^{\prime}\}$, $B=B_{0}\cup B_{1}$ and order $B$ as follows:
$1<\overline{1}<2<\overline{2}<\cdots<n<\overline{n}<1^{\prime}<2^{\prime}<\cdots<m^{\prime}.$
An orthosymplectic tableau $T$ of shape $\lambda$ is a filling of Young diagram of $\lambda$ with entries from $B$ satisfying the following conditions:
  • the portion $S$ of $T$ that contains entries only from $B_{0}$ is symplectic.
  • the skew tableau formed by removing S from T is strictly increasing across the rows from left to right and weakly increasing down columns from top to bottom.
The weight of such a tableau is given by
$\operatorname{wt}(T)=\prod_{i=1}^{n}x_{i}^{n_{\overline{i}}(T)-n_{i}(T)}\prod_{j=1}^{m}y_{j}^{n_{i^{\prime}}(T)},$
where $n_{\alpha}$ is the number of occurrences of $\alpha$, for $\alpha$ being any of the entries given in 2.5. Denote the set of orthosymplectic tableaux of shape $\lambda$ filled with entries $1\dots m^{\prime}$ by $\operatorname{OSPT}_{n/m}({\lambda})$.
Example 2.6.
The following is an orthosymplectic tableau $T$ of shape $\lambda=(3,3,2,1)$ and $\operatorname{wt}(T)=x_{1}x_{4}^{-1}y_{1}y_{2}y_{3}^{3}$.
{ytableau}$\ytableau\overline{1}&23^{\prime}\\ \overline{2}1^{\prime}3^{\prime}\\ 43^{\prime}\\ 2^{\prime}$
The orthosymplectic Schur function indexed by a partition $\lambda$ is given by
$\operatorname{spo}_{\lambda}(X/Y)\coloneqq\sum_{T\,\in\,\operatorname{OSPT}_{n/m}({\lambda})}\operatorname{wt}(T).$
Note that $\operatorname{spo}_{\lambda}(X/Y)$ is zero unless $\lambda_{n+1}\leq m$. Also, $\operatorname{spo}_{\lambda}(X/\emptyset)=\mathrm{sp}_{\lambda}(X)$. Moreover,
$\operatorname{spo}_{\lambda}(X/Y)=\sum_{\mu}\mathrm{sp}_{\mu}(X)s_{\lambda^{\prime}/\mu^{\prime}}(Y).$
Example 2.7.
Let $n=1$, $m=2$ and consider the partition $\lambda=(2)$. Then we have the following orthosymplectic tableaux in
[…]
Therefore,
$\operatorname{spo}_{(2)}(x_{1}/y_{1},y_{2})=x_{1}^{2}+\bar{x}_{1}^{2}+1+y_{1}(x_{1}+\bar{x}_{1})+y_{2}(x_{1}+\bar{x}_{1})+y_{1}y_{2}.$
The Jacobi–Trudi type identity for the orthosymplectic Schur functions is given by
(2.13) $\operatorname{spo}_{\lambda}(X/Y)=\det\left(J_{\lambda_{i}-i+j}+J_{\lambda_{i}-i-j+2}\right)_{n\times n},$
where $J_{r}=\sum_{l=0}^{r}h_{l}(X,\bar{X})e_{r-l}(Y)$.
Note that the orthosymplectic Schur polynomial $\operatorname{spo}_{\lambda}(X/Y)$ is nonzero if and only if $\lambda_{n+1}\leq m$. The following remark gives an expression for $\operatorname{spo}_{\lambda}(X/y)$ if $\lambda_{n+1}\leq 1$.
Remark 2.8.
[…]
$\operatorname{spo}_{\lambda}(X/y)=\frac{y^{\ell-n}}{D}\det\left(x_{i}^{\lambda_{j}+n-j+1}-\bar{x}_{i}^{\lambda_{j}+n-j+1}+y(x_{i}^{\lambda_{j}+n-j}-\bar{x}_{i}^{\lambda_{j}+n-j})\right),$
where $D=\prod_{i=1}^{n}(x_{i}-\bar{x}_{i})\prod_{1\leq i<j\leq n}(x_{i}+\bar{x}_{i}-x_{j}-\bar{x}_{j})$.
Now we give a relation between odd symplectic characters defined by Proctor [Pro88] and orthosymplectic Schur functions after defining odd symplectic characters.
Definition 2.9.
Let $\lambda$ be a partition of length at most $n$. A $(n-1,1)$-symplectic tableau of shape $\lambda$ is a filling of the Young diagram of $\lambda$ with entries from
$1<\overline{1}<2<\overline{2}<\cdots<n-1<\overline{n-1}<n$
satisfying the following three conditions:
  • the entries in each row are weakly increasing;
  • the entries in each column are strictly increasing;
  • the entries of the $i$’th row are greater than or equal to $i$
We denote by $\operatorname{SPT}^{(n-1,1)}(\lambda)$ the set of $(n-1,1)$-symplectic tableaux of shape $\lambda$. Given a $(n-1,1)$-symplectic tableau $T$, we define
$\operatorname{wt}(T)=x_{n}^{\alpha_{n}(T)}\prod_{i=1}^{n-1}x_{i}^{\alpha_{i}(T)-\alpha_{\overline{i}}(T)},$
where $\alpha_{i}(T)$, $\alpha_{\overline{i}}(T)$ for $i\in[n-1]$, and $\alpha_{n}(T)$ are number of occurences of $i$, $\overline{i}$ and $n$ respectively.
The odd symplectic character indexed by a partition of length at most $n$ is given by
$\mathrm{sp}^{(n-1,1)}(x_{1},\dots,x_{n-1}|x_{n})=\sum_{T\in\operatorname{SPT}^{(n-1,1)}(\lambda)}\operatorname{wt}(T).$
Okada [Oka20] proved the following determinantal formula for odd symplectic characters.
Theorem 2.10 ([Oka20, Theorem 1.1]).
For a partition of length at most $n$, let $A_{\lambda}=(a_{ij})$ be a $n\times n$ matrix such that
$a_{ij}=\begin{cases}(x_{i}^{\lambda_{j}+n-j+1}-\bar{x}_{i}^{\lambda_{j}+n-j+1})-y^{-1}(x_{i}^{\lambda_{j}+n-j}-\bar{x}_{i}^{\lambda_{j}+n-j})&1\leq i\leq n-1,\\ y^{\lambda_{j}+n-j}&i=n.\end{cases}$
Then
$\mathrm{sp}^{(n-1,1)}_{\lambda}(x_{1},\dots,x_{n-1}|x_{n})=\frac{\det A_{\lambda}}{\det A_{\emptyset}},$
where
$\det A_{\emptyset}=\prod_{i=1}^{n-1}(x_{i}-x_{i}^{-1})\prod_{1\leq i<j\leq n}(x_{i}+\bar{x}_{i}-x_{j}-\bar{x}_{j}).$
Comparing the determinantal formulas in 1.4 and 2.10, we get the following result.
Corollary 2.11.
Suppose $\lambda$ is a partition of length at most $n$. Then
$\mathrm{sp}^{(n-1,1)}_{\lambda}(x_{1},\dots,x_{n-1}|x_{n})=\operatorname{spo}_{\lambda}(x_{1},\dots,x_{n-1},x_{n}^{-1}/-x_{n}^{-1}).$
3. Determinant formula for orthosymplectic Schur functions
In this section, we give a proof of 1.2. We first prove an equivalent statement of 1.2 given in 3.3. Before that, let us recall the following lemmas.
Lemma 3.1 ([FH91, Lemma A.54]).
Let $\zeta_{j}(p)=x_{j}^{p}-\bar{x}_{j}^{p}$, $1\leq j\leq n$. Then for any integer $l\geq 0$, $\zeta_{j}(l)$ is the product of the $1\times n$, $n\times n$ and $n\times 1$ matrices
$\left(\begin{array}[]{cccc}\overline{H}_{l-n}&\overline{H}_{l-n+1}+\overline{H}_{l-n-1}&\dots&\overline{H}_{l-1}+\overline{H}_{l-2n+1}\end{array}\right)\left(\begin{array}[]{c}(-1)^{q-p}e_{q-p}\end{array}\right)\left(\begin{array}[]{c}\zeta_{j}(n)\\ \zeta_{j}(n-1)\\ \vdots\\ \zeta_{j}(1)\end{array}\right),$
where $\overline{H}_{r}=h_{r}(X,\bar{X})$ for all $r\in\mathbb{Z}$.
Lemma 3.2 ([Mac15, Chapter I.1, Equation (1.7)]).
Let $\lambda$ be a partition such that $\ell(\lambda)\leq n_{1}$ and $\lambda_{1}\leq n_{2}$. Then the $n_{1}+n_{2}$ numbers
$\lambda_{i}+n_{1}-i\quad(1\leq i\leq n_{1}),\qquad\qquad n_{1}-1+j-\lambda_{j}^{\prime}\quad(1\leq j\leq n_{2})$
are a permutation of $\{0,1,\dots,n_{1}+n_{2}-1\}$.
Theorem 3.3.
Let $\lambda$ be a partition of length at most $n$ and $k=\min\{j|\lambda_{j}+n+1-j\leq m\}.$ Then the orthosymplectic Schur function $\operatorname{spo}_{\lambda}(X/Y)$ is given by
(3.1) $\operatorname{spo}_{\lambda}(X/Y)=\frac{(-1)^{mn-n+k-1}}{D}\det\left(\begin{array}[]{c|c}R&X_{\lambda}\\[-8.5359pt] \\ \hline\cr\\[-8.5359pt] Y_{\lambda}&0\end{array}\right),$
where
$R=\left((-1)^{m-j}(x_{i}^{j}-\bar{x}_{i}^{j})\right)_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq m\end{subarray}},\quad Y_{\lambda}=\left(h_{\lambda^{\prime}_{i}-n-i+j}(Y)\right)_{\begin{subarray}{c}1\leq i\leq m-n+k-1\\ 1\leq j\leq m\end{subarray}},$
$X_{\lambda}=\left({x_{i}^{\lambda_{j}+n-m-j+1}}{\prod_{q=1}^{m}(x_{i}+y_{q})}-{\bar{x}_{i}^{\lambda_{j}+n-m-j+1}}{\prod_{q=1}^{m}(\bar{x}_{i}+y_{q})}\right)_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq k-1\end{subarray}},$
and
$D={\prod_{i=1}^{n}(x_{i}-\bar{x}_{i})\prod_{1\leq i<j\leq n}(x_{i}+\bar{x}_{i}-x_{j}-\bar{x}_{j})}.$
Proof.
By Jacobi–Trudi type identity for orthosymplectic Schur functions (2.13), we have
(3.2) $\operatorname{spo}_{\lambda}(X/Y)=\det\left(J_{\lambda_{i}-i+j}+J_{\lambda_{i}-i-j+2}\right)_{1\leq i,j\leq n},$
where $J_{r}=\displaystyle\sum_{p=0}^{r}h_{p}(X,\bar{X})e_{r-p}(Y)=\displaystyle\sum_{p=0}^{r}h_{r-p}(X,\bar{X})e_{p}(Y)$ for all $r\geq 0$. Otherwise $J_{r}=0$. Since $e_{p}(Y)=0$ for $p>m$ and $h_{r-p}(X,\bar{X})=0$, for $p>r$,
(3.3) $J_{r}=\displaystyle\sum_{p=0}^{m}h_{r-p}(X,\bar{X})e_{p}(Y).,\quad r\in\mathbb{Z}.$
Substituting the expression for $J_{r}$ from (3.3) in (3.2), we have
$\operatorname{spo}_{\lambda}(X/Y)=\det\left(\sum_{p=0}^{m}(\overline{H}_{\lambda_{i}-i+j-p}+\overline{H}_{\lambda_{i}-i-j+2-p})e_{p}(Y)\right)_{n\times n},$
where $\overline{H}_{r}=h_{r}(X,\bar{X})$ for all $r\in\mathbb{Z}$. Let $k=\text{min}\{j|\lambda_{j}+n+1-j\leq m\}.$ Now applying 3.1, the required orthosymplectic Schur function is given by
$\frac{1}{D}\det\left(\begin{array}[]{c}\left(\displaystyle\sum_{p=0}^{m}\left(x_{j}^{\lambda_{i}+n-i+1-p}-\bar{x}_{j}^{\lambda_{i}+n-i+1-p}\right)e_{p}(Y)\right)_{1\leq i\leq k-1}\\[-8.5359pt] \\ \hline\cr\\[-8.5359pt] \left(\displaystyle\sum_{p=0}^{\lambda_{i}+n-i}\left(x_{j}^{\lambda_{i}+n-i+1-p}-\bar{x}_{j}^{\lambda_{i}+n-i+1-p}\right)e_{p}(Y)\right)_{k\leq i\leq n}\end{array}\right)_{1\leq j\leq n},$
where
$D=\displaystyle\prod_{i=1}^{n}(x_{i}-\bar{x}_{i})\prod_{1\leq i<j\leq n}(x_{i}+\bar{x}_{i}-x_{j}-\bar{x}_{j}).$
Note that
(3.4) $\sum_{p=0}^{m}x_{i}^{-p}e_{p}(Y)=\prod_{j=1}^{m}(1+\bar{x}_{i}y_{j}),\qquad\sum_{p=0}^{m}x_{i}^{p}e_{p}(Y)=\prod_{i=1}^{m}(1+x_{i}y_{j}).$
Consider the permutation $\sigma=(k,\dots,n,1,\dots,k-1)$, in one line notation in $S_{n}$. Then, the number of inversions of $\sigma$ is $(k-1)(n-k+1)$ and it can be seen that
(3.5) $\operatorname{spo}_{\lambda}(X/Y)=\frac{\operatorname{sgn}(\sigma)}{D}\det\left(\begin{array}[]{c|c}(A_{i,j})_{\begin{subarray}{c}1\leq i\leq n\\ k\leq j\leq n\end{subarray}}&(B_{i,j})_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq k-1\end{subarray}}\end{array}\right),$
where
$A_{i,j}=\sum_{p=0}^{\lambda_{j}+n-j}\left(x_{i}^{\lambda_{j}+n-j+1-p}-\bar{x}_{i}^{\lambda_{j}+n-j+1-p}\right)e_{p}(Y)$
and
$B_{i,j}=\sum_{p=0}^{m}\left(x_{i}^{\lambda_{j}+n-j+1-p}-\bar{x}_{i}^{\lambda_{j}+n-j+1-p}\right)e_{p}(Y)\\ =x_{i}^{\lambda_{j}+n-j+1}\left(\prod_{q=1}^{m}(1+\bar{x}_{i}y_{q})\right)-\bar{x}_{i}^{\lambda_{j}+n-j+1}\left(\prod_{q=1}^{m}(1+x_{i}y_{q})\right),$
where the last equality uses the relations given in (3.4). Let $\Lambda\coloneqq\{\lambda_{j}+n-j+1:k\leq j\leq n\}$ and $\Lambda^{\prime}\coloneqq\{n+i-\lambda_{i}^{\prime}:1\leq i\leq m-n+k-1\}$. Note that $\{\Lambda,\Lambda^{\prime}\}$ is a set partition of $\{1,2,\dots,m\}$ by 3.2. So, we can rewrite (3.5) in the following manner.
$\operatorname{spo}_{\lambda}(X/Y)=\frac{(-1)^{b}}{D}\det\left(\begin{array}[]{c|c|c}((-1)^{m-j}(x_{i}^{j}-\bar{x}_{i}^{j}))_{\begin{subarray}{c}1\leq i\leq n\\ j\,\in\,\Lambda^{\prime}\end{subarray}}&(A_{i,j})_{\begin{subarray}{c}1\leq i\leq n\\ k\leq j\leq n\end{subarray}}&(B_{i,j})_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq k-1\end{subarray}}\\ &&\\ \hline\cr&\\ (h_{-i+j}(Y))_{\begin{subarray}{c}i\,\in\,\Lambda^{\prime}\\ j\,\in\,\Lambda^{\prime}\end{subarray}}&{\text{\huge 0}}&{\text{\huge 0}}\end{array}\right),$
where $b=(k-1)(n-k+1)+n(m-n+k-1)=mn-(n-k+1)^{2}$. Observe that
$(A_{i,j})_{\begin{subarray}{c}1\leq i\leq n\\ k\leq j\leq n\end{subarray}}=\left(\sum_{p=0}^{j-1}(x_{i}^{j-p}-\bar{x}_{i}^{j-p})e_{p}(Y)\right)_{\begin{subarray}{c}1\leq i\leq n\\ j\,\in\,\Lambda\end{subarray}}$
So, the orthosymplectic Schur function $\operatorname{spo}_{\lambda}(X/Y)$ is given by
(3.6) $\frac{(-1)^{b}}{D}\det\left(\begin{array}[]{c|c|c}\left((-1)^{m-j}\phi_{i}(j)\right)_{\begin{subarray}{c}1\leq i\leq n\\ j\,\in\,\Lambda^{\prime}\end{subarray}}&\left(\sum_{p=0}^{j-1}\phi_{i}(j-p)e_{p}(Y)\right)_{\begin{subarray}{c}1\leq i\leq n\\ j\,\in\,\Lambda\end{subarray}}&(B_{i,j})_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq k-1\end{subarray}}\\ &&\\ \hline\cr&\\ (h_{-i+j}(Y))_{\begin{subarray}{c}i\,\in\,\Lambda^{\prime}\\ j\,\in\,\Lambda^{\prime}\end{subarray}}&{\text{\huge 0}}&{\text{\huge 0}}\end{array}\right),$
where $\phi_{i}(j)=x_{i}^{j}-\bar{x}_{i}^{j}$. Applying $C_{\lambda_{n}+1}\rightarrow C_{\lambda_{n}+1}-\sum_{p=1}^{\lambda_{n}}(-1)^{m-p}e_{\lambda_{n}+1-p}(Y)C_{p}$ and using
$-\sum_{p=1}^{\lambda_{n}}(-1)^{m-p}e_{\lambda_{n}+1-p}(Y)h_{-i+p}(Y)=(-1)^{m-\lambda_{n}+1}h_{-i+\lambda_{n}+1}(Y),$
the function in (3.6) is the same as
$\frac{(-1)^{b}}{D}\\ \times\left|\begin{array}[]{c|c|c|c}\left((-1)^{m-j}\phi_{i}(j)\right)_{\begin{subarray}{c}1\leq i\leq n\\ j\,\in\,\Lambda^{\prime}\end{subarray}}&\left(\displaystyle\sum_{p=0}^{j-1}\psi_{i}(p)e_{p}(Y)\right)_{\begin{subarray}{c}1\leq i\leq n\\ j\,\in\,\Lambda\end{subarray}}&(\phi_{i}(j))_{\begin{subarray}{c}1\leq i\leq n\\ j=\lambda_{n}+1\end{subarray}}&(B_{i,j})_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq k-1\end{subarray}}\\ &&&\\ \hline\cr&&\\ (h_{-i+j}(Y))_{\begin{subarray}{c}i\,\in\,\Lambda^{\prime}\\ j\,\in\,\Lambda^{\prime}\end{subarray}}&{\text{\huge 0}}&((-1)^{m-j}h_{-i+j}(Y))_{\begin{subarray}{c}i\in\Lambda^{\prime}\\ j=\lambda_{n}+1\end{subarray}}&{\text{\huge 0}}\end{array}\right|,$
where $\psi_{i}(p)=\phi_{i}(j-p)$. Successively applying $C_{j}\rightarrow C_{j}-\sum_{p=1}^{j-1}(-1)^{m-p}e_{j-p}C_{p}$, for $j\in\Lambda$, we have
$\operatorname{spo}_{\lambda}(X/Y)=\frac{(-1)^{b}}{D}\det\left(\begin{array}[]{c|c|c}\left((-1)^{m-j}\phi(j)\right)_{\begin{subarray}{c}1\leq i\leq n\\ j\,\in\,\Lambda^{\prime}\end{subarray}}&(\phi(j))_{\begin{subarray}{c}1\leq i\leq n\\ j\,\in\,\Lambda\end{subarray}}&(B_{i,j})_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq k-1\end{subarray}}\\ &&\\ \hline\cr&\\ (h_{-i+j}(Y))_{\begin{subarray}{c}i\,\in\,\Lambda^{\prime}\\ j\,\in\,\Lambda^{\prime}\end{subarray}}&((-1)^{m-j}h_{-i+j}(Y))_{\begin{subarray}{c}i\,\in\,\Lambda^{\prime}\\ j\,\in\,\Lambda\end{subarray}}&{\text{\huge 0}}\end{array}\right).$
Rearranging the columns, we have
$\operatorname{spo}_{\lambda}(X/Y)=\frac{(-1)^{a}}{D}\det\left(\begin{array}[]{c|c}\left((-1)^{m-j}(x_{i}^{j}-\bar{x}_{i}^{j})\right)_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq m\end{subarray}}&(B_{i,j})_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq k-1\end{subarray}}\\[17.50002pt] \hline\cr\\ \left(h_{\lambda^{\prime}_{i}-n-i+j}(Y)\right)_{\begin{subarray}{c}1\leq i\leq m-n+k-1\\ 1\leq j\leq m\end{subarray}}&{\text{\huge 0}}\end{array}\right),$
where $a=mn-n+k-1$. This completes the proof. ∎
Now we prove that the expressions for $\operatorname{spo}_{\lambda}(X/Y)$ in 1.2 and 3.3 are the same.
Proof of 1.2.
Assume that (1.1) holds true. Let $a=mn-n+k-1$. Then the orthosymplectic Schur function $\operatorname{spo}_{\lambda}(X/Y)$ is given by
(3.7) $\frac{(-1)^{a}}{D_{s}}\det\left(\begin{array}[]{c|c}\left(\displaystyle\frac{x_{i}}{(x_{i}+y_{j})\prod_{q=1}^{m}(\bar{x}_{i}+y_{q})}-\displaystyle\frac{\bar{x}_{i}}{(\bar{x}_{i}+y_{j})\prod_{q=1}^{m}(x_{i}+y_{q})}\right)_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq m\end{subarray}}&X_{\lambda}\\[17.50002pt] \hline\cr\\ \left(y_{j}^{\lambda^{\prime}_{i}+m-n-i}\right)_{\begin{subarray}{c}1\leq i\leq m-n+k-1\\ 1\leq j\leq m\end{subarray}}&{\text{\huge 0}}\end{array}\right),$
where
$X_{\lambda}=\left(\frac{x_{i}^{\lambda_{j}+n-m-j+1}}{\prod_{q=1}^{m}(\bar{x}_{i}+y_{q})}-\frac{\bar{x}_{i}^{\lambda_{j}+n-m-j+1}}{\prod_{q=1}^{m}(x_{i}+y_{q})}\right)_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq k-1\end{subarray}}$
and
$D_{s}=\frac{\prod_{i=1}^{n}(x_{i}-\bar{x}_{i})\prod_{1\leq i<j\leq n}(x_{i}+\bar{x}_{i}-x_{j}-\bar{x}_{j})\prod_{1\leq i<j\leq m}(y_{i}-y_{j})}{\prod_{i=1}^{n}\prod_{j=1}^{m}(x_{i}+y_{j})(\bar{x}_{i}+y_{j})}.$
Applying $R_{i}\rightarrow\prod_{q=1}^{m}(x_{i}+y_{q})(\bar{x}_{i}+y_{q})R_{i}$, for $i\in[1,n]$ and cancelling the term $\prod_{i=1}^{n}\prod_{j=1}^{m}(x_{i}+y_{j})(\bar{x}_{i}+y_{j})$ from the denominator of $D_{s}$, (3.7) can be written as
$\frac{(-1)^{a}}{D_{s}^{1}}\det\left(\begin{array}[]{c|c}\left(x_{i}\prod_{\begin{subarray}{c}q=1\\ q\neq j\end{subarray}}^{m}(x_{i}+y_{q})-\bar{x}_{i}\prod_{\begin{subarray}{c}q=1\\ q\neq j\end{subarray}}^{m}(\bar{x}_{i}+y_{q})\right)_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq m\end{subarray}}&Z_{\lambda}\\[17.50002pt] \hline\cr\\ \left(y_{j}^{\lambda^{\prime}_{i}+m-n-i}\right)_{\begin{subarray}{c}1\leq i\leq m-n+k-1\\ 1\leq j\leq m\end{subarray}}&{\text{\huge 0}}\end{array}\right),$
where
$D_{s}^{1}=\prod_{i=1}^{n}(x_{i}-\bar{x}_{i})\prod_{1\leq i<j\leq n}(x_{i}+\bar{x}_{i}-x_{j}-\bar{x}_{j})\prod_{1\leq i<j\leq m}(y_{i}-y_{j})$
and
$Z_{\lambda}=\left({x_{i}^{\lambda_{j}+n-m-j+1}}{\prod_{q=1}^{m}(x_{i}+y_{q})}-{\bar{x}_{i}^{\lambda_{j}+n-m-j+1}}{\prod_{q=1}^{m}(\bar{x}_{i}+y_{q})}\right)_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq k-1\end{subarray}},$
Let
$M^{k}_{i,j}=x_{i}\prod_{\begin{subarray}{c}q=1\\ q\neq j\end{subarray}}^{k}(x_{i}+y_{q})-\bar{x}_{i}\prod_{\begin{subarray}{c}q=1\\ q\neq j\end{subarray}}^{k}(\bar{x}_{i}+y_{q}).$
Applying $C_{j}\rightarrow C_{j}-C_{m}$, $j\in[1,m-1]$ and then using $y_{j}^{p}-y_{m}^{p}=(y_{j}-y_{m})h_{p-1}(y_{j},y_{m})$, the orthosymplectic Schur function $\operatorname{spo}_{\lambda}(X/Y)$ is given by
$\frac{(-1)^{a}}{D_{s}^{2}}\det\left(\begin{array}[]{c|c|c}\left(-M^{m-1}_{i,j}\right)_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq m-1\end{subarray}}&\left(M^{m}_{i,j}\right)_{1\leq i\leq n}&Z_{\lambda}\\ &\\ \hline\cr&\\ \left(h_{\lambda^{\prime}_{i}+m-n-i-1}(y_{j},y_{m})\right)_{\begin{subarray}{c}1\leq i\leq m-n+k-1\\ 1\leq j\leq m-1\end{subarray}}&\left(h_{\lambda^{\prime}_{i}+m-n-i}(y_{m})\right)_{1\leq i\leq m-n+k-1}&{\text{\huge 0}}\end{array}\right),$
where
$D_{s}^{2}=\prod_{i=1}^{n}(x_{i}-\bar{x}_{i})\prod_{1\leq i<j\leq n}(x_{i}+\bar{x}_{i}-x_{j}-\bar{x}_{j})\prod_{1\leq i<j\leq m-1}(y_{i}-y_{j}).$
Now applying $C_{j}\rightarrow C_{j}-C_{s}$, for $j\in[1,s-1]$ and then using $h_{p}(y_{j},y_{s+1},y_{s+2},\dots,y_{m})-h_{p}(y_{s},y_{s+1},y_{s+2},\dots,y_{m})=(y_{j}-y_{s})h_{p-1}(y_{j},y_{s},y_{s+1},y_{s+2},\dots,y_{m})$, successively for $s=m-1,m-2,\dots,2$, we have
$\operatorname{spo}_{\lambda}(X/Y)=\frac{(-1)^{a}}{D}\det\left(\begin{array}[]{c|c}\left((-1)^{m-j}(x_{i}\prod_{q=1}^{j-1}(x_{i}+y_{q})-\bar{x}_{i}\prod_{q=1}^{j-1}(\bar{x}_{i}+y_{q}))\right)_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq m\end{subarray}}&Z_{\lambda}\\[17.50002pt] \hline\cr\\ \left(h_{\lambda^{\prime}_{i}-n-i+j}(y_{j},\dots,y_{m})\right)_{\begin{subarray}{c}1\leq i\leq m-n+k-1\\ 1\leq j\leq m\end{subarray}}&{\text{\huge 0}}\end{array}\right),$
where
$D=\prod_{i=1}^{n}(x_{i}-\bar{x}_{i})\prod_{1\leq i<j\leq n}(x_{i}+\bar{x}_{i}-x_{j}-\bar{x}_{j}).$
Applying $C_{j}\rightarrow C_{j}+y_{j-1}C_{j-1}$, for $j\in[2,m]$ and using $h_{n}(y_{j-1},\dots,y_{m})=h_{n}(y_{j},\dots,y_{m})+y_{j-1}h_{n-1}(y_{j-1},\dots,y_{m})$, $\operatorname{spo}_{\lambda}(X/Y)$ is given by
$\frac{(-1)^{a}}{D}\det\left(\begin{array}[]{c|c|c}\left((-1)^{m-1}(x_{i}-\bar{x}_{i})\right)_{1\leq i\leq n}&\left(N_{i,j}^{2}\right)_{\begin{subarray}{c}1\leq i\leq n\\ 2\leq j\leq m\end{subarray}}&Z_{\lambda}\\ &&\\ \hline\cr&&\\ \left(h_{\lambda^{\prime}_{i}-n-i+1}(Y)\right)_{1\leq i\leq m-n+k-1}&\left(h_{\lambda^{\prime}_{i}-n-i+j}(y_{j-1},\dots,y_{m})\right)_{\begin{subarray}{c}1\leq i\leq m-n+k-1\\ 2\leq j\leq m\end{subarray}}&{\text{\huge 0}}\end{array}\right),$
where
$N_{i,j}^{s}=(-1)^{m-j}\left(x_{i}^{s}\prod_{q=1}^{j-s}(x_{i}+y_{q})-\bar{x}_{i}^{s}\prod_{q=1}^{j-s}(\bar{x}_{i}+y_{q})\right).$
Now again applying $C_{j}\rightarrow C_{j}+y_{j-1}C_{j-1}$, for $j\in[3,m]$, we see that the orthosymplectic Schur function is given by
$\frac{(-1)^{a}}{D}\det\left(\begin{array}[]{c|c|c}\left((-1)^{m-j}(x_{i}^{j}-\bar{x}_{i}^{j})\right)_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq 2\end{subarray}}&\left(N_{i,j}^{3}\right)_{\begin{subarray}{c}1\leq i\leq n\\ 3\leq j\leq m\end{subarray}}&Z_{\lambda}\\ &&\\ \hline\cr&&\\ \left(h_{\lambda^{\prime}_{i}-n-i+j}(Y)\right)_{\begin{subarray}{c}1\leq i\leq m-n+k-1\\ 1\leq j\leq 2\end{subarray}}&\left(h_{\lambda^{\prime}_{i}-n-i+j}(y_{j-2},\dots,y_{m})\right)_{\begin{subarray}{c}1\leq i\leq m-n+k-1\\ 2\leq j\leq m\end{subarray}}&{\text{\huge 0}}\end{array}\right).$
Proceeding in a similar fashion, we have
(3.8) $\operatorname{spo}_{\lambda}(X/Y)=\frac{(-1)^{a}}{D}\det\left(\begin{array}[]{c|c}\left((-1)^{m-j}(x_{i}^{j}-\bar{x}_{i}^{j})\right)_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq m\end{subarray}}&Z_{\lambda}\\[17.50002pt] \hline\cr\\ \left(h_{\lambda^{\prime}_{i}-n-i+j}(Y)\right)_{\begin{subarray}{c}1\leq i\leq m-n+k-1\\ 1\leq j\leq m\end{subarray}}&{\text{\huge 0}}\end{array}\right).$
Observe that the expression in (3.8) is the same as the expression in (3.1). This completes the proof. ∎
4. A generalization of Brent–Krattenthaler–Warnaar’s identity
In this section, we prove a generalization of the following odd symplectic character identity due to Brent, Krattenthaler and Warnaar, which was found in the study of discrete Mehta-type integrals [BKW16]. Recently, Okada gave its linear-algebraic proof [Oka20].
Theorem 4.1 ([Oka20, Theorem 1.2]).
Let $m$ and $n$ be positive integers with $n\leq m$ and $r$ a non-negative integer. Then
$\sum_{\lambda}z^{-r}\mathrm{sp}^{(n,1)}_{\lambda}(X|z)\,\mathrm{sp}^{(m,1)}_{(r^{m-n})\cup\lambda}(Y|z)=\mathrm{sp}_{(\underbrace{r,\dots,r}_{m+n+1})}(X,Y,z),$
where $\lambda$ runs over partitions of length at most $n+1$ such that $\lambda_{1}\leq r$, and $(r^{m-n})\cup\lambda$ stands for the partition $(\underbrace{r,\dots,r}_{m-n},\lambda_{1},\dots,\lambda_{n+1})$.
We write some lemmas before presenting the generalization (see 4.7). Recall the Cauchy–Binet formula. Given an $m\times n$ matrix $X=(x_{i,j})_{{1\leq i\leq m,1\leq j\leq n}}$ and a subset $B\subset[n]=\{1,\dots,n\}$ of column indices, let $X[B]$ be the submatrix of $X$ obtained by picking up colmns indexed by elements of $B$.
Lemma 4.2.
For two $m\times n$ matrices $X$ and $Y$, we have
$\sum_{B}\det X[B]\det Y[B]=\det(XY^{\prime}),$
where the sum runs over $m$-elements subset of $[n]$.
Lemma 4.3 ([Oka20, Lemma 3.1(b)]).
Let $\lambda$ be a partition of length at most $n$ and $\lambda_{1}\leq r$. Then
$(x_{1}\cdots x_{n})^{r}\mathrm{sp}_{\lambda}(X)$
is a polynomial in $x_{1},\dots,x_{n}$, and we have
$\begin{split}[(x_{1}\cdots x_{n})^{r}\mathrm{sp}_{\lambda}(X)]&|_{x_{1}=0}\\ =&\begin{cases}(x_{2}\cdots x_{n})^{r}\mathrm{sp}_{(\lambda_{2},\dots,\lambda_{n})}(x_{2},\dots,x_{n})&\text{ if }\lambda_{1}=r,\\ 0&\text{ otherwise,}\end{cases}\end{split}$
where $F|_{x_{1}=0}$ means that we substitute $x_{1}=0$ in $F$.
We need the following generalization of 4.3 in the proof of 4.7.
Lemma 4.4.
Let $\lambda$ be a partition of length at most $n$ and $\lambda_{1}\leq r$. Then
$(x_{1}\cdots x_{n})^{r}\operatorname{spo}_{\lambda}(X/z)$
is a polynomial in $x_{1},\dots,x_{n}$, and we have
$\begin{split}[(x_{1}\cdots x_{n})^{r}\operatorname{spo}_{\lambda}(X/z)]&|_{x_{1}=0}\\ =&\begin{cases}(x_{2}\cdots x_{n})^{r}\operatorname{spo}_{(\lambda_{2},\dots,\lambda_{n})}(x_{2},\dots,x_{n}/z)&\text{ if }\lambda_{1}=r,\\ 0&\text{ otherwise,}\end{cases}\end{split}$
where $F|_{x_{1}=0}$ means that we substitute $x_{1}=0$ in $F$.
Proof.
By 1.4, we have
$\operatorname{spo}_{\lambda}(X/z)=\frac{1}{D}\det\left(x_{i}^{\lambda_{j}+n-j+1}-\bar{x}_{i}^{\lambda_{j}+n-j+1}+z\left(x_{i}^{\lambda_{j}+n-j}-\bar{x}_{i}^{\lambda_{j}+n-j}\right)\right),$
where $D=\prod_{i=1}^{n}(x_{i}-\bar{x}_{i})\prod_{1\leq i<j\leq n}(x_{i}+\bar{x}_{i}-x_{j}-\bar{x}_{j})$. Let $A_{\lambda}(X)\coloneqq(a_{i,j})_{1\leq i,j\leq n}$ such that
$\begin{split}a_{i,j}=x_{i}^{r+n}\left(x_{i}^{\lambda_{j}+n-j+1}-\bar{x}_{i}^{\lambda_{j}+n-j+1}+z\left(x_{i}^{\lambda_{j}+n-j}-\bar{x}_{i}^{\lambda_{j}+n-j}\right)\right)&\\ =x_{i}^{\lambda_{j}+r+2n-j+1}-x_{i}^{r-\lambda_{j}+j-1}+&z\left(x_{i}^{\lambda_{j}+r+2n-j}-x_{i}^{r-\lambda_{j}+j}\right).\end{split}$
Then
$(x_{1}\cdots x_{n})^{r}\operatorname{spo}_{\lambda}(X/z)=\frac{\det A_{\lambda}(X)}{x_{1}^{n}\cdots x_{n}^{n}}\times\frac{1}{D}.$
Since $\lambda_{j}+r+2n-j+1>\lambda_{j}+r+2n-j>r-\lambda_{j}+j>r-\lambda_{j}+j-1\geq 0,$ $(x_{1}\cdots x_{n})^{r}\operatorname{spo}_{\lambda}(X/z)$ is a polynomial in $x_{1},\dots,x_{n}$. Also, note that $r-\lambda_{j}+j-1=0$ if and only if $j=1$ and $\lambda_{1}=r$. Therefore,
$a_{1,j}|_{x_{1}=0}=\begin{cases}-1&\lambda_{1}=r\text{ and }j=1\\ 0&\text{ otherwise.}\end{cases}$
And
$(\det A_{\lambda}(X))|_{x_{1}=0}=\begin{cases}-\det\left(a_{i+1,j+1}\right)_{1\leq i,j\leq n-1}&\lambda_{1}=r,\\ 0&\text{ otherwise.}\end{cases}$
Observe that
$\det\left(a_{i+1,j+1}\right)_{1\leq i,j\leq n-1}=(x_{2}\cdots x_{n})\det A_{(\lambda_{2},\dots,\lambda_{n})}(x_{2},\dots,x_{n}).$
If $\lambda_{1}=r$, then we have
$\begin{split}[(x_{1}\cdots x_{n})^{r}\operatorname{spo}_{\lambda}(X/z)]|_{x_{1}=0}&=\frac{-(x_{2}\cdots x_{n})\det A_{(\lambda_{2},\dots,\lambda_{n})}(x_{2},\dots,x_{n})}{-(x_{2}\cdots x_{n})^{n}\prod_{i=2}^{n}(x_{i}-\bar{x}_{i})\prod_{2\leq i<j\leq n}(x_{i}+\bar{x}_{i}-x_{j}-\bar{x}_{j})}\\ &=(x_{2}\cdots x_{n})^{r}\operatorname{spo}_{(\lambda_{2},\dots,\lambda_{n})}(x_{2},\dots,x_{n}/z).\end{split}$
This completes the proof. ∎
Lemma 4.5 ([Oka20, Lemma 3.2]).
Let $p(x,y,z,a,b)$ be a rational function in $x$, $y$, $z$, $a$ and $b$ given by
$\begin{split}p(x,y,&z,a,b)=\\ &\frac{(1-xz)(1-yz)}{1-xy}-a\frac{(x-z)(1-yz)}{x-y}+b\frac{(1-xz)(y-z)}{x-y}-ab\frac{(x-z)(y-z)}{1-xy}.\end{split}$
Let $\mathbf{x}=(x_{1},\dots,x_{n}),$ $\mathbf{y}=(y_{1},\dots,y_{n}),$ $\mathbf{a}=(a_{1},\dots,a_{n}),$ $\mathbf{b}=(b_{1},\dots,b_{n}),$ $z$ and $c$ be determinates. We define an $(n+1)\times(n+1)$ matrix $C=C(\mathbf{x},\mathbf{y},z;\mathbf{a},\mathbf{b},c)=(C_{i,j})_{1\leq i,j\leq n+1}$ by putting
$C_{i,j}=\begin{cases}p(x_{i},y_{j},z,a_{i},b_{j})&1\leq i,j\leq n,\\ 1-a_{i}&1\leq i\leq n,j=n+1,\\ 1-b_{j}&i=n+1,1\leq j\leq n,\\ \displaystyle\frac{1-c}{1-z^{2}}&i=j=n+1,\end{cases}$
and a $(2n+1)\times(2n+1)$ matrix $V=V(\mathbf{x},\mathbf{y},z;\mathbf{a},\mathbf{b},c)=(V_{i,j})_{1\leq i,j\leq 2n+1}$ by putting
$V_{i,j}=\begin{cases}x_{i}^{j-1}-a_{i}x_{i}^{2n+1-j}&1\leq i\leq n,\\ y_{i-n}^{j-1}-b_{i-n}y_{i-n}^{2n+1-j}&n+1\leq i\leq 2n,\\ z^{j-1}-cz^{2n+1-j}&i=2n+1,\end{cases}$
for $11\leq j\leq 2n+$ 1. Then
(4.1) $\det C=\frac{(-1)^{n}}{\displaystyle(1-z^{2})\prod_{i=1}^{n}\prod_{j=1}^{n}(x_{i}-y_{j})(1-x_{i}y_{j})}\det V.$
Lemma 4.6.
Let $q(x,y,z,a,b)$ be a rational function in $x$, $y$, $z$, $a$ and $b$ given by
(4.2) $\begin{split}q(x,y,&z,a,b)=\\ &\frac{(1+xz)(1+yz)}{1-xy}-a\frac{(x+z)(1+yz)}{x-y}+b\frac{(1+xz)(y+z)}{x-y}-ab\frac{(x+z)(y+z)}{1-xy}.\end{split}$
Then
$\det\left(q(x_{i},y_{j},z,a_{i},b_{j})\right)_{1\leq i,j\leq n}=\frac{(-1)^{n}}{\displaystyle\prod_{i=1}^{n}\prod_{j=1}^{n}(x_{i}-y_{j})(1-x_{i}y_{j})}\det W,$
where $W=(W_{i,j})_{1\leq i,j\leq 2n+1}$ such that
(4.3) $W_{i,j}=\begin{cases}x_{i}^{j-1}-a_{i}x_{i}^{2n+1-j}&1\leq i\leq n\\ y_{i-n}^{j-1}-b_{i-n}y_{i-n}^{2n+1-j}&n+1\leq i\leq 2n\\ (-z)^{2n+1-j}&i=2n+1,\end{cases}$
for $11\leq j\leq 2n+$ 1.
Proof.
Recall the notions of $p(x,y,z,a,b)$, $(n+1)\times(n+1)$ matrix $C$ and $(2n+1)\times(2n+1)$ matrix $V$ from 4.5. Observe that
$q(x,y,z,a,b)=p(x,y,-z,a,b)$
and
$\det\left(q(x_{i},y_{j},-z,a_{i},b_{j})\right)_{1\leq i,j\leq n}=\frac{(1-z^{2})}{-c}\left(\det C-\det C|_{c=0}\right).$
Then by (4.1), we have
$\det\left(q(x_{i},y_{j},-z,a_{i},b_{j})\right)_{1\leq i,j\leq n}=\frac{(-1)^{n}}{\displaystyle-c\prod_{i=1}^{n}\prod_{j=1}^{n}(x_{i}-y_{j})(1-x_{i}y_{j})}\left(\det V-\det V|_{c=0}\right)$
Therefore,
$\det\left(q(x_{i},y_{j},z,a_{i},b_{j})\right)_{1\leq i,j\leq n}=\frac{(-1)^{n}}{\displaystyle\prod_{i=1}^{n}\prod_{j=1}^{n}(x_{i}-y_{j})(1-x_{i}y_{j})}\det W,$
where $W$ is given in (4.3). This completes the proof. ∎
Theorem 4.7.
Let $m$ and $n$ be positive integers with $n\leq m$, and $r$ a nonnegative integer. Then we have
(4.4) $\begin{split}\sum_{\lambda}z^{r}\operatorname{spo}_{\lambda}(X/z)\operatorname{spo}_{(r^{m-n})\cup\lambda}&(Y/z)\\ =\sum_{j=1}^{m+n+1}&z^{r+m+n+1-j}\mathrm{sp}_{\big{(}\underbrace{r,\dots,r}_{j-1},\underbrace{r-1,\dots,r-1}_{m+n+1-j}\big{)}}(X,Y),\end{split}$
where $\lambda$ runs over partitions of length at most $n$ such that $\lambda_{1}\leq r$.
Proof.
Assume $m=n$. Let $M=r+n$ and consider $n\times M$ matrices $X=(X_{i,p})_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq p\leq M\end{subarray}}$ and $Y=(Y_{i,p})_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq p\leq M\end{subarray}}$ such that
$\begin{cases}X_{i,p}&=\quad x_{i}^{p}-\bar{x}_{i}^{p}+z(x_{i}^{p-1}-\bar{x}_{i}^{p-1}),\\ Y_{i,p}&=\quad y_{i}^{p}-\bar{y}_{i}^{p}+z(y_{i}^{p-1}-\bar{y}_{i}^{p-1}),\end{cases}$
for $1\leq i\leq n$, $1\leq p\leq M$. Then from a direct but lengthy computation, one can see that the $(i,j)$’th entry of $XY^{\prime}$ is given by
$\bar{x}_{i}^{M}\bar{y}_{j}^{M}q(x_{i},y_{j},z,x_{i}^{2M},y_{j}^{2M}),$
where
$q(x,y,z,a,b)=\\ \frac{(1+xz)(1+yz)}{1-xy}-a\frac{(x+z)(1+yz)}{x-y}+b\frac{(1+xz)(y+z)}{x-y}-ab\frac{(x+z)(y+z)}{1-xy}.$
Therefore,
(4.5) $\det XY^{\prime}=(\bar{x}_{1}\cdots\bar{x}_{n})^{M}(\bar{y}_{1}\cdots\bar{y}_{n})^{M}\det\left(q(x_{i},y_{j},z,x_{i}^{2M},y_{j}^{2M})\right).$
By 4.6, we have
(4.6) $\det\left(q(x_{i},y_{j},z,x_{i}^{2M},y_{j}^{2M})\right)=\frac{(-1)^{n}}{\displaystyle\prod_{i=1}^{n}\prod_{j=1}^{n}(x_{i}-y_{j})(1-x_{i}y_{j})}\det W^{(M)},$
where
$\left(W^{(M)}\right)_{i,j}=\begin{cases}x_{i}^{j-1}-x_{i}^{2M+2n+1-j}&1\leq i\leq n\\ y_{i-n}^{j-1}-y_{i-n}^{2M+2n+1-j}&n+1\leq i\leq 2n\\ (-z)^{2n+1-j}&i=2n+1,\end{cases}$
for $11\leq j\leq 2n+$ 1. Suppose $I_{n}(\lambda)=(\lambda_{1}+n,\dots,\lambda_{n}+1)$. Then by 1.4 and the Cauchy–Binet formula in 4.2, the left hand side of (4.4) is given by
(4.7) $\begin{split}\sum_{\lambda}z^{r}\operatorname{spo}_{\lambda}(X/z)\operatorname{spo}_{\lambda}(Y/z)=&\displaystyle\sum_{\lambda}\frac{z^{r}}{D_{x}D_{y}}\det X[I_{n}(\lambda)]\,{\det Y[I_{n}(\lambda)]}\\ =\frac{z^{r}}{D_{x}D_{y}}\det(XY^{\prime})=&\frac{(-1)^{n}z^{r}(\bar{x}_{1}\cdots\bar{x}_{n})^{M}(\bar{y}_{1}\cdots\bar{y}_{n})^{M}}{D_{x}D_{y}\displaystyle\prod_{i=1}^{n}\prod_{j=1}^{n}(x_{i}-y_{j})(1-x_{i}y_{j})}\times\det W^{(M)},\end{split}$
where the last equality uses (4.5) and (4.6) and $D_{x}=\prod_{i=1}^{n}(x_{i}-\bar{x}_{i})\prod_{1\leq i<j\leq n}(x_{i}+\bar{x}_{i}-x_{j}-\bar{x}_{j})$, $D_{y}=\prod_{i=1}^{n}(y_{i}-\bar{y}_{i})\prod_{1\leq i<j\leq n}(y_{i}+\bar{y}_{i}-y_{j}-\bar{y}_{j})$. Simplifying (4.7), we have
$\sum_{\lambda}z^{r}\operatorname{spo}_{\lambda}(X/z)\operatorname{spo}_{\lambda}(Y/z)=\frac{(-1)^{r+n}\prod_{i=1}^{n}x_{i}^{n}y_{i}^{n}}{D_{x}D_{y}\displaystyle\prod_{i=1}^{n}\prod_{j=1}^{n}(x_{i}-y_{j})(1-x_{i}y_{j})}\det W_{0},$
where
$(W_{0})_{i,j}=\begin{cases}x_{i}^{M+n+1-j}-\bar{x}_{i}^{M+n+1-j}&1\leq i\leq n,\\ y_{i}^{M+n+1-j}-\bar{y}_{i}^{M+n+1-j}&n+1\leq i\leq 2n,\\ (-z)^{M+n+1-j}&i=2n+1,\end{cases}$
for $11\leq j\leq 2n+$ 1. Note that
$\frac{\displaystyle\prod_{i=1}^{n}\prod_{j=1}^{n}(x_{i}-y_{j})(1-x_{i}y_{j})}{\prod_{i=1}^{n}x_{i}^{n}y_{i}^{n}}=\prod_{i=1}^{n}\prod_{j=1}^{n}(y_{j}+\bar{y}_{j}-x_{i}-\bar{x}_{i})=(-1)^{n^{2}}\prod_{i=1}^{n}\prod_{j=1}^{n}(x_{i}+\bar{x}_{i}-y_{j}-\bar{y}_{j}).$
Therefore,
$\begin{split}\sum_{\lambda}z^{r}\operatorname{spo}_{\lambda}(X/z)\operatorname{spo}_{\lambda}(Y/z)=\frac{(-1)^{r+n^{2}+n}}{D_{x}D_{y}\prod_{i=1}^{n}\prod_{j=1}^{n}(x_{i}+\bar{x}_{i}-y_{j}-\bar{y}_{j})}&\det W_{0}\\ =\sum_{j=1}^{2n+1}z^{r+2n+1-j}\mathrm{sp}_{\big{(}\underbrace{r,\dots,r}_{j-1},\underbrace{r-1,\dots,r-1}_{2n+1-j}\big{)}}&(X,Y),\end{split}$
where the last equality uses (2.11). This completes the proof in the case $m=n$. Now we prove the general case by downward induction on $n$. We assume (4.4). Multiplying both sides of (4.4) by $(x_{1}\cdots x_{n}y_{1}\cdots y_{m})^{r}$ and then substituting $x_{1}=0$, by 4.3 and 4.4, we have
$\begin{split}\sum_{\lambda}z^{r}(x_{2}\cdots x_{n})^{r}\operatorname{spo}_{(\lambda_{2},\dots,\lambda_{n})}(x_{2},\dots,x_{n}/z)(y_{1}\cdots y_{m})^{r}\operatorname{spo}_{(r^{m-n})\cup\lambda}(Y&/z)\\ =\sum_{j=2}^{m+n+1}z^{r+m+n+1-j}(x_{2}\cdots x_{n}y_{1}\cdots y_{m})^{r}\mathrm{sp}_{\big{(}\underbrace{r,\dots,r}_{j-2},\underbrace{r-1,\dots,r-1}_{m+n+1-j}\big{)}}&(x_{2},\dots,x_{n},y_{1},\dots,y_{m}),\end{split}$
where the summation on the left hand side is over partitions $\lambda$ of length at most $n$ such that $\lambda_{1}=r$. Since $(r^{m-n})\cup\lambda=(r^{m-n+1})\cup(\lambda_{2},\dots,\lambda_{n})$ for such a partition, we have
$\begin{split}\sum_{\mu}z^{r}\operatorname{spo}_{\mu}(x_{2},\dots,x_{n}/z)\operatorname{spo}_{(r^{m-n+1})\,\cup\,\mu}(Y&/z)\\ =\sum_{j=1}^{m+n}z^{r+m+n-j}\mathrm{sp}_{\big{(}\underbrace{r,\dots,r}_{j-1},\underbrace{r-1,\dots,r-1}_{m+n-j}\big{)}}&(x_{2},\dots,x_{n},y_{1},\dots,y_{m}),\end{split}$
where $\mu$ runs over all partitions of length at most $n-1$ such that $\mu_{1}\leq r$. This completes the proof. ∎
Acknowledgements
We thank A. Ayyer and A. Stokke for very helpful discussions.

### v2

@@
 These are supersymmetric functions in two sets of commuting variables indexed by integer partitions.
-It is well known that hook Schur polynomials form the basis for the ring of supersymmetric functions.
+It is well known that hook Schur polynomials form a basis for the ring of supersymmetric functions.
 For background on the representation theory of Lie superalgebras, see [CW12].
@@
 Stembridge [Ste85] gave a different characterization for these polynomials and showed that they form a basis for the ring of supersymmetric functions.
-Moens and Jeugt [MvdJ03] proved the following determinantal formula for these polynomials using the representation theory of general linear superalgebra $\mathfrak{gl}(m/n)$.
+Moens and Jeugt [MvdJ03] proved the following determinantal formula for these polynomials using the representation theory of the general linear superalgebra $\mathfrak{gl}(m/n)$.
 Theorem 1.1 ([MvdJ03, Equation 1.17]).
-Let $\lambda$ be a partition such that $\lambda_{n+1}\leq m$ and $k=\min\{i|\lambda_{i}+n+1-i\leq m\}.$ Then the hook Schur polynomial $\operatorname{hs}_{\lambda}(X/Y)$ is given by
-$\operatorname{hs}_{\lambda}(X/Y)=\frac{(-1)^{mn-n+k-1}}{D}\det\left(\begin{array}[]{c|c}\left(\frac{1}{x_{i}+y_{j}}\right)_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq m\end{subarray}}&\left(x_{i}^{\lambda_{j}+n-m-j}\right)_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq k-1\end{subarray}}\\ \\ \hline\cr\\ \left(y_{j}^{\lambda^{\prime}_{i}+m-n-i}\right)_{\begin{subarray}{c}1\leq i\leq m-n+k-1\\ 1\leq j\leq m\end{subarray}}&{\text{\huge 0}}\end{array}\right),$
+Let $\lambda$ be a partition such that $\lambda_{n+1}\leq m$ and $k=\min\{i\mid\lambda_{i}+n+1-i\leq m\}.$ The hook Schur polynomial $\operatorname{hs}_{\lambda}(X;Y)$ is given by
+$\operatorname{hs}_{\lambda}(X;Y)=\frac{(-1)^{mn-n+k-1}}{D}\det\left(\begin{array}[]{c|c}\left(\frac{1}{x_{i}+y_{j}}\right)_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq m\end{subarray}}&\left(x_{i}^{\lambda_{j}+n-m-j}\right)_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq k-1\end{subarray}}\\ \\ \hline\cr\\ \left(y_{j}^{\lambda^{\prime}_{i}+m-n-i}\right)_{\begin{subarray}{c}1\leq i\leq m-n+k-1\\ 1\leq j\leq m\end{subarray}}&{\text{\huge 0}}\end{array}\right),$
 where
@@
 Theorem 1.2.
-Let $\lambda$ be a partition of length at most $n$ and $k=\min\{j|\lambda_{j}+n+1-j\leq m\}.$ Then the orthosymplectic Schur function $\operatorname{spo}_{\lambda}(X/Y)$ is given by
-(1.1) $\operatorname{spo}_{\lambda}(X/Y)=\frac{(-1)^{mn-n+k-1}}{D}\det\left(\begin{array}[]{c|c}R&X_{\lambda}\\[-8.5359pt] \\ \hline\cr\\[-8.5359pt] Y_{\lambda}&0\end{array}\right)$
+Let $\lambda$ be a partition of length at most $n$ and $k=\min\{j\mid\lambda_{j}+n+1-j\leq m\}.$ The orthosymplectic Schur function $\operatorname{spo}_{\lambda}(X;Y)$ is given by
+(1.1) $\operatorname{spo}_{\lambda}(X;Y)=\frac{(-1)^{mn-n+k-1}}{D}\det\left(\begin{array}[]{c|c}R&X_{\lambda}\\[-8.5359pt] \\ \hline\cr\\[-8.5359pt] Y_{\lambda}&0\end{array}\right)$
 where
@@
 Let $n=1$, $m=2$ and consider the partition $\lambda=(2)$. Then $k=2$ and by (1.1), we have
-$\operatorname{spo}_{\lambda}(x_{1}/y_{1},y_{2})=\frac{(x_{1}+y_{1})(x_{1}+y_{2})(\bar{x}_{1}+y_{1})(\bar{x}_{1}+y_{2})}{(x_{1}-\bar{x}_{1})(y_{1}-y_{2})}\\ \times\left(\frac{x_{1}}{(\bar{x}_{1}+y_{1})(\bar{x}_{1}+y_{2})}-\frac{\bar{x}_{1}}{(x_{1}+y_{1})(x_{1}+y_{2})}\right)\times\det\left(\begin{array}[]{cc}y_{1}&y_{2}\\ 1&1\end{array}\right).$
+$\operatorname{spo}_{\lambda}(x_{1};y_{1},y_{2})=\frac{(x_{1}+y_{1})(x_{1}+y_{2})(\bar{x}_{1}+y_{1})(\bar{x}_{1}+y_{2})}{(x_{1}-\bar{x}_{1})(y_{1}-y_{2})}\\ \times\left(\frac{x_{1}}{(\bar{x}_{1}+y_{1})(\bar{x}_{1}+y_{2})}-\frac{\bar{x}_{1}}{(x_{1}+y_{1})(x_{1}+y_{2})}\right)\times\det\left(\begin{array}[]{cc}y_{1}&y_{2}\\ 1&1\end{array}\right).$
 It turns out to be equal to
-$x_{1}^{2}+\bar{x}_{1}^{2}+1+y_{1}(x_{1}+\bar{x}_{1})+y_{2}(x_{1}+\bar{x}_{1})+y_{1}y_{2}.$
+(1.2) $x_{1}^{2}+\bar{x}_{1}^{2}+1+y_{1}(x_{1}+\bar{x}_{1})+y_{2}(x_{1}+\bar{x}_{1})+y_{1}y_{2}.$
+As the expression for $\operatorname{spo}_{(2)}(x_{1};y_{1},y_{2})$ in (1.2) is the same as the one given in (2.12), 1.2 is verified in this instance.
 Corollary 1.4.
 Let $\lambda$ be a partition of length at most $n$. Then
-$\operatorname{spo}_{\lambda}(x_{1},x_{2},\dots,x_{n}/y)=\frac{1}{D}\det\left(x_{i}^{\lambda_{j}+n-j+1}-\bar{x}_{i}^{\lambda_{j}+n-j+1}+y\left(x_{i}^{\lambda_{j}+n-j}-\bar{x}_{i}^{\lambda_{j}+n-j}\right)\right),$
+$\operatorname{spo}_{\lambda}(x_{1},x_{2},\dots,x_{n};y)=\frac{1}{D}\det\left(x_{i}^{\lambda_{j}+n-j+1}-\bar{x}_{i}^{\lambda_{j}+n-j+1}+y\left(x_{i}^{\lambda_{j}+n-j}-\bar{x}_{i}^{\lambda_{j}+n-j}\right)\right),$
 where $D=\prod_{i=1}^{n}(x_{i}-\bar{x}_{i})\prod_{1\leq i<j\leq n}(x_{i}+\bar{x}_{i}-x_{j}-\bar{x}_{j})$.
@@
 The Jacobi–Trudi type formula was given by Balantekin and Bars [BB81].
-It was proved using a different approach by Benkart, Shader and Ram [BSR98] and recently, Stokke and Visentin [SV16] constructed the formula using lattice paths.
-Orthosymplectic Cauchy identities and Pieri rule for orthosymplectic Schur polynomials were proved in [PPS22] and  [Sto18] respectively.
-We also show that the odd symplectic characters introduced by Proctor [Pro88] are the same as the orthosymplectic Schur polynomials with some specialized indeterminates (see 2.11).
+It was proved using a different approach by Benkart, Shader and Ram [BSR98] and recently, Stokke and Visentin [SV16] proved the formula using lattice paths.
+Orthosymplectic Cauchy identities and a Pieri rule for orthosymplectic Schur polynomials were proved in [PPS22] and  [Sto18] respectively.
+We also show that the odd symplectic characters introduced by Proctor [Pro88] are the same as the orthosymplectic Schur polynomials with some specialized indeterminates (see 2.12).
 Lastly, we give a generalization of an odd symplectic character identity due to Brent, Krattenthaler and Warnaar, which was found in the study of discrete Mehta-type integrals (see  [BKW16, BKW15, Oka20]) in Section 4.
 2. Preliminaries
-Recall that a partition $\lambda=(\lambda_{1},\lambda_{2},\dots,\lambda_{\ell})$ is a sequence of weakly decreasing nonnegative integers.
-Any entry of the partition is called a part.
+Recall that a partition $\lambda=(\lambda_{1},\lambda_{2},\dots)$ is a sequence of weakly decreasing nonnegative integers containing only finitely many non-zero terms.
 The length of a partition $\lambda$, which is the number of positive parts, is denoted by $\ell(\lambda)$.
-A partition $\lambda$ can be represented pictorially as a Young diagram, whose $i$’th row contains left-justified boxes.
+A partition $\lambda$ can be represented pictorially as a Young diagram, whose $i^{\text{th}}$ row contains $\lambda_{i}$ left-justified boxes.
 We will use the so-called English notation where the first row is on top.
 The conjugate of a partition $\lambda$, denoted $\lambda^{\prime}$, is the partition whose Young diagram is obtained by transposing the Young diagram of $\lambda$.
+For example, Figure 1 shows the Young diagrams of $\lambda=(3,2,2)$ and its conjugate $\lambda^{\prime}=(3,3,1)$.
+[…]
+3,2,2
+[…]
+3,3,1
+Figure 1. The Young diagram of $\lambda=(3,2,2)$ and $\lambda^{\prime}=(3,3,1)$
 Throughout, we fix $n$ and $m$ to be positive integers.
@@
 For a partition $\lambda=(\lambda_{1},\ldots,\lambda_{n})$ of length at most $n$, the Schur polynomial is given by the following determinantal formula:
-(2.3) $s_{\lambda}(X)=\frac{\displaystyle\det_{1\leq i,j\leq n}\Bigl{(}x_{i}^{\lambda_{j}+n-j}\Bigr{)}}{\displaystyle\det_{1\leq i,j\leq n}\Bigl{(}x_{i}^{n-j}\Bigr{)}}.$
-The expression in (2.3) is essentially the Weyl character formula. The following Jacobi–Trudi identity expresses the Schur polynomial $s_{\lambda}(X)$ in terms of complete symmetric functions,
-(2.4) $s_{\lambda}(X)=\det\left(h_{\lambda_{i}-i+j}(X)\right)_{1\leq i,j\leq r},$
-where $r$ is any integer such that $r\geq\ell(\lambda)$.
-Recall that there is a combinatorial formula for $s_{\lambda}(X)$ as a sum of monomials summed over all column-strict Young tableaux of shape $\lambda$.
-A column-strict Young tableau of shape $\lambda$ is a filling of $\lambda$ with entries from the set $[n]$ such that
-  • the entries increase weakly along rows
-  • the entries increase strictly along columns.
-For partitions $\lambda$ and $\mu$ such that $\mu\subset\lambda\,(\mu_{i}\leq\lambda_{i}$, for all $i\geq 1)$, the skew shape $\lambda/\mu$ is the set theoretic difference $\lambda/\mu$ and the skew Schur function $s_{\lambda/\mu}(X)$ is given by
-(2.5) $s_{\lambda/\mu}(X)=\sum_{T}\operatorname{wt}(T),\qquad\operatorname{wt}(T)\coloneqq\prod_{i\geq 1}x_{i}^{c_{i}(T)},$
-where the sum is taken over all column-strict tableaux of shape $\lambda/\mu$ and $c_{i}(T)$, $i\in[n]$ is the number of occurrences of $i$ in $T$.
-Otherwise $s_{\lambda/\mu}(X)=0$.
-The following Jacobi–Trudi formula expresses $s_{\lambda/\mu}(X)$ in terms of complete symmetric functions.
-(2.6) $s_{\lambda/\mu}(X)=\det\left(h_{\lambda_{i}-\mu_{j}-i+j}(X)\right)_{1\leq i,j\leq r},$
+(2.3) $\mathrm{s}_{\lambda}(X)=\frac{\displaystyle\det\Bigl{(}x_{i}^{\lambda_{j}+n-j}\Bigr{)}_{1\leq i,j\leq n}}{\displaystyle\det\Bigl{(}x_{i}^{n-j}\Bigr{)}_{1\leq i,j\leq n}}.$
+The expression in (2.3) is essentially the Weyl character formula.
+For partitions $\lambda$ and $\mu$ such that $\mu\subset\lambda\,(\mu_{i}\leq\lambda_{i}$, for all $i\geq 1)$, the skew shape $\lambda/\mu$ is the set theoretic difference where $\lambda$ and $\mu$ are regarded as Young diagrams, i.e., collections of boxes.
+Recall that a semistandard Young tableau of shape $\lambda/\mu$ is a filling of $\lambda/\mu$ such that
+  • the entries increase weakly across rows,
+  • the entries increase strictly down columns.
+The skew Schur function $\mathrm{s}_{\lambda/\mu}(X)$ is given by
+(2.4) $\mathrm{s}_{\lambda/\mu}(X)=\sum_{T}\operatorname{wt}(T),\qquad\operatorname{wt}(T)\coloneqq\prod_{i=1}^{n}x_{i}^{c_{i}(T)},$
+where the sum is taken over all semistandard Young tableaux of shape $\lambda/\mu$ with entries from the set $[n]$ and $c_{i}(T)$, $i\in[n]$ is the number of occurrences of $i$ in $T$.
+If $\mu=\emptyset$, then $\mathrm{s}_{\lambda/\mu}(X)=\mathrm{s}_{\lambda}(X)$.
+The following Jacobi–Trudi formula expresses $\mathrm{s}_{\lambda/\mu}(X)$ in terms of complete symmetric functions:
+(2.5) $\mathrm{s}_{\lambda/\mu}(X)=\det\left(h_{\lambda_{i}-\mu_{j}-i+j}(X)\right)_{1\leq i,j\leq r},$
 where $r$ is any integer such that $r\geq\ell(\lambda)$.
@@
 A polynomial $f(X,Y)$ in this ring is doubly symmetric if it is symmetric in both $(x_{1},\dots,x_{n})$ and $(y_{1},\dots,y_{m})$.
-Moreover, if substituting $x_{n}=t$ and $y_{m}=-t$ results in an expression independent of $t$, then we call $f(X,Y)$ a supersymmetric function.
+Moreover, if substituting $x_{n}=t$ and $y_{m}=-t$ results in an element in $\mathbb{Z}[x_{1},\dots,x_{n-1},y_{1},\dots,y_{m-1}]$, then we call $f(X,Y)$ a supersymmetric function.
 The elementary supersymmetric function and the complete supersymmetric function indexed by $\lambda=(\lambda_{1},\lambda_{2},\dots,\lambda_{n})$ are:
-(2.7) $E_{\lambda}(X/Y)=\prod_{i=1}^{n}E_{\lambda_{i}}(X/Y),$
+(2.6) $E_{\lambda}(X;Y)=\prod_{i=1}^{n}E_{\lambda_{i}}(X;Y),$
 where
-$E_{r}(X/Y)=\begin{cases}\displaystyle\sum_{j=0}^{r}e_{j}(X)h_{r-j}(Y)&r\geq 1,\\ \qquad\qquad 1&r=0,\end{cases}$
+$E_{r}(X;Y)=\begin{cases}\displaystyle\sum_{j=0}^{r}e_{j}(X)h_{r-j}(Y)&r\geq 1,\\ \qquad\qquad 1&r=0,\end{cases}$
 and
-(2.8) $H_{\lambda}(X/Y)=\prod_{i=1}^{n}H_{\lambda_{i}}(X/Y),$
+(2.7) $H_{\lambda}(X;Y)=\prod_{i=1}^{n}H_{\lambda_{i}}(X;Y),$
 where
-$H_{r}(X/Y)=\begin{cases}\displaystyle\sum_{j=0}^{r}h_{j}(X)e_{r-j}(Y)&r\geq 1,\\ \qquad\qquad 1&r=0.\end{cases}$
-It is convenient to define $E_{r}(X/Y)$ and $H_{r}(X/Y)$ to be zero for $r<0$.
+$H_{r}(X;Y)=\begin{cases}\displaystyle\sum_{j=0}^{r}h_{j}(X)e_{r-j}(Y)&r\geq 1,\\ \qquad\qquad 1&r=0.\end{cases}$
+It is convenient to define $E_{r}(X;Y)$ and $H_{r}(X;Y)$ to be zero for $r<0$.
 Definition 2.1.
@@
 is a filling of the shape with these entries satisfying the following conditions:
-  • the entries increase weakly along rows and columns
-  • the unprimed entries strictly increase along columns
-  • the primed entries strictly increase along rows
+  • the entries increase weakly across rows and down columns,
+  • the unprimed entries strictly increase down columns,
+  • the primed entries strictly increase across rows.
 The weight of such a supertableau is given by
-$\operatorname{wt}(T)\coloneqq\prod_{i=1}^{n}x_{i}^{n_{i}(T)}\prod_{j=1^{\prime}}^{m^{\prime}}y_{j}^{n_{j}(T)},$
+$\operatorname{wt}(T)\coloneqq\prod_{i=1}^{n}x_{i}^{n_{i}(T)}{\prod_{j=1}^{m}y_{j}^{n_{j^{\prime}}(T)}},$
 where $n_{\alpha}(T)$, for $\alpha$ being any of the entries given in 2.1, is the number of occurrences of $\alpha$ in $T$.
 Denote the set of supertableaux of skew shape $\lambda$ with entries given in 2.1 by $\operatorname{SSYT}_{n/m}(\lambda)$.
-For an integer partition $\lambda$, the hook Schur (or supersymmetric Schur) polynomial, denoted $\operatorname{hs}_{\lambda}(X/Y)$ is given by
-(2.9) $\operatorname{hs}_{\lambda/\mu}(X/Y)\coloneqq\sum_{T\,\in\,\operatorname{SSYT}_{n/m}(\lambda)}\operatorname{wt}(T).$
+For an integer partition $\lambda$, the hook Schur (or supersymmetric Schur) polynomial, denoted $\operatorname{hs}_{\lambda}(X;Y)$ is given by
+(2.8) $\operatorname{hs}_{\lambda/\mu}(X;Y)\coloneqq\sum_{T\,\in\,\operatorname{SSYT}_{n/m}(\lambda)}\operatorname{wt}(T).$
 Example 2.2.
@@
 Therefore,
-$\operatorname{hs}_{(2,1)}(x_{1},x_{2}/y_{1})=x_{1}^{2}x_{2}+x_{1}x_{2}^{2}+x_{1}^{2}y_{1}+2x_{1}x_{2}y_{1}+x_{2}^{2}y_{1}+x_{1}y_{1}^{2}+x_{2}y_{1}^{2}.$
-The Jacobi–Trudi formula expresses $\operatorname{hs}_{\lambda}(X/Y)$ in terms of the complete supersymmetric functions [PT92].
-(2.10) $\operatorname{hs}_{\lambda}(X/Y)=\det\left(H_{\lambda_{i}-i+j}(X/Y)\right)_{1\leq i,j\leq n}.$
-Note that the hook Schur polynomial $\operatorname{hs}_{\lambda}(X/Y)$ is nonzero if and only if $\lambda_{n+1}\leq m$.
+$\operatorname{hs}_{(2,1)}(x_{1},x_{2};y_{1})=x_{1}^{2}x_{2}+x_{1}x_{2}^{2}+x_{1}^{2}y_{1}+2x_{1}x_{2}y_{1}+x_{2}^{2}y_{1}+x_{1}y_{1}^{2}+x_{2}y_{1}^{2}.$
+The Jacobi–Trudi formula expresses $\operatorname{hs}_{\lambda}(X;Y)$ in terms of the complete supersymmetric functions [PT92].
+(2.9) $\operatorname{hs}_{\lambda}(X;Y)=\det\left(H_{\lambda_{i}-i+j}(X;Y)\right)_{1\leq i,j\leq n}.$
+Note that the hook Schur polynomial $\operatorname{hs}_{\lambda}(X;Y)$ is nonzero if and only if $\lambda_{n+1}\leq m$.
 Now we consider orthosymplectic Schur functions after defining symplectic Schur functions.
 Definition 2.3.
-A symplectic tableau $T$ of shape $\lambda$ with entries
-$1<\overline{1}<2<\overline{2}<\cdots<n<\overline{n}$
-is a filling of the shape with these entries satisfying the following conditions:
-  • the entries increase weakly along rows and strictly along columns
-[…]
+A symplectic tableau $T$ of shape $\lambda$ is a filling of the shape with entries from $1<\overline{1}<2<\overline{2}<\cdots<n<\overline{n}$ satisfying the following conditions:
+  • the entries increase weakly across rows and strictly down columns,
+  • the entries in the $i^{\text{th}}$ row must be greater than or equal to $i$.
 The weight of such a tableau is given by
-$\operatorname{wt}(T)=\prod_{i=1}^{n}x_{i}^{n_{\overline{i}}(T)-n_{i}(T)},$
-where $n_{i}(T)$ and $n_{\overline{i}}(T)$ is the number of occurrences of $i$ and $\overline{i}$ in $T$ respectively.
-Denote the set of symplectic tableaux of shape $\lambda$ filled with entries given in 2.3 by $\operatorname{SPT}_{n}(\lambda)$.
+$\operatorname{wt}(T)={\prod_{i=1}^{n}x_{i}^{n_{i}(T)-n_{\overline{i}}(T)},}$
+where $n_{i}(T)$ and $n_{\overline{i}}(T)$ are the number of occurrences of $i$ and $\overline{i}$ in $T$ respectively.
+Denote the set of symplectic tableaux of shape $\lambda$ filled with entries given in 2.3 by $\operatorname{SP}_{n}(\lambda)$.
 Example 2.4.
-The following is a symplectic tableau $T$ of shape $\lambda=(3,2,2)$ with $\operatorname{wt}(T)=x_{2}$.
+The following is a symplectic tableau $T$ of shape $\lambda=(3,2,2)$ with $\operatorname{wt}(T)={x_{2}^{-1}}$.
 {ytableau}$\ytableau 1&\overline{1}2\\ \overline{2}\overline{2}\\ 4\overline{4}$
 The symplectic Schur function is defined as
-$\mathrm{sp}_{\lambda}(X)\coloneqq\sum_{T\in\operatorname{SPT}_{n}(\lambda)}\operatorname{wt}(T),$
-where the sum runs over the set of tableaux in $\operatorname{SPT}_{n}(\lambda)$.
+$\mathrm{sp}_{\lambda}(X)\coloneqq\sum_{T\in\operatorname{SP}_{n}(\lambda)}\operatorname{wt}(T),$
+where the sum runs over the set of tableaux in $\operatorname{SP}_{n}(\lambda)$.
 The symplectic Schur function can be expressed using the Weyl character formula in the following way.
@@
 Then
-(2.11) $\mathrm{sp}_{\lambda}(X)=\frac{\displaystyle\det_{1\leq i,j\leq n}\Bigl{(}x_{i}^{\lambda_{j}+n-j+1}-\bar{x}_{i}^{\lambda_{j}+n-j+1}\Bigr{)}}{\displaystyle\det_{1\leq i,j\leq n}\Bigl{(}x_{i}^{n-j+1}-\bar{x}_{i}^{n-j+1}\Bigr{)}},$
+(2.10) $\mathrm{sp}_{\lambda}(X)=\frac{\displaystyle\det\Bigr{(}x_{i}^{\lambda_{j}+n-j+1}-\bar{x}_{i}^{\lambda_{j}+n-j+1}\Bigr{)}_{1\leq i,j\leq n}}{\displaystyle\det\Bigr{(}x_{i}^{n-j+1}-\bar{x}_{i}^{n-j+1}\Bigr{)}_{1\leq i,j\leq n}},$
 and the denominator here is
-(2.12) $\det_{1\leq i,j\leq n}\Bigl{(}x_{i}^{n-j+1}-\bar{x}_{i}^{n-j+1}\Bigr{)}=\prod_{i=1}^{n}(x_{i}-\bar{x}_{i})\,\prod_{1\leq i<j\leq n}(x_{i}+\bar{x}_{i}-x_{j}-\bar{x}_{j}).$
+(2.11) $\det\Bigr{(}x_{i}^{n-j+1}-\bar{x}_{i}^{n-j+1}\Bigr{)}_{1\leq i,j\leq n}=\prod_{i=1}^{n}(x_{i}-\bar{x}_{i})\,\prod_{1\leq i<j\leq n}(x_{i}+\bar{x}_{i}-x_{j}-\bar{x}_{j}).$
 Definition 2.5.
@@
 $1<\overline{1}<2<\overline{2}<\cdots<n<\overline{n}<1^{\prime}<2^{\prime}<\cdots<m^{\prime}.$
-An orthosymplectic tableau $T$ of shape $\lambda$ is a filling of Young diagram of $\lambda$ with entries from $B$ satisfying the following conditions:
-  • the portion $S$ of $T$ that contains entries only from $B_{0}$ is symplectic.
+An orthosymplectic tableau $T$ of shape $\lambda$ is a filling of the Young diagram of $\lambda$ with entries from $B$ satisfying the following conditions:
+  • the portion $S$ of $T$ that contains entries only from $B_{0}$ is symplectic,
   • the skew tableau formed by removing S from T is strictly increasing across the rows from left to right and weakly increasing down columns from top to bottom.
 The weight of such a tableau is given by
-$\operatorname{wt}(T)=\prod_{i=1}^{n}x_{i}^{n_{\overline{i}}(T)-n_{i}(T)}\prod_{j=1}^{m}y_{j}^{n_{i^{\prime}}(T)},$
-where $n_{\alpha}$ is the number of occurrences of $\alpha$, for $\alpha$ being any of the entries given in 2.5.
-Denote the set of orthosymplectic tableaux of shape $\lambda$ filled with entries $1\dots m^{\prime}$ by $\operatorname{OSPT}_{n/m}({\lambda})$.
+$\operatorname{wt}(T)={\prod_{i=1}^{n}x_{i}^{n_{i}(T)-n_{\overline{i}}(T)}\prod_{j=1}^{m}y_{j}^{n_{i^{\prime}}(T)}},$
+where $n_{\alpha}$ is the number of occurrences of $\alpha$, for $\alpha$ being any of the entries in $B$.
+Denote the set of orthosymplectic tableaux of shape $\lambda$ filled with entries in $B$ by $\operatorname{OSP}_{n/m}({\lambda})$.
 Example 2.6.
-The following is an orthosymplectic tableau $T$ of shape $\lambda=(3,3,2,1)$ and $\operatorname{wt}(T)=x_{1}x_{4}^{-1}y_{1}y_{2}y_{3}^{3}$.
+The following is an orthosymplectic tableau $T$ of shape $\lambda=(3,3,2,1)$ and $\operatorname{wt}(T)=x_{1}^{-1}x_{4}y_{1}y_{2}y_{3}^{3}$.
 {ytableau}$\ytableau\overline{1}&23^{\prime}\\ \overline{2}1^{\prime}3^{\prime}\\ 43^{\prime}\\ 2^{\prime}$
 The orthosymplectic Schur function indexed by a partition $\lambda$ is given by
-$\operatorname{spo}_{\lambda}(X/Y)\coloneqq\sum_{T\,\in\,\operatorname{OSPT}_{n/m}({\lambda})}\operatorname{wt}(T).$
-Note that $\operatorname{spo}_{\lambda}(X/Y)$ is zero unless $\lambda_{n+1}\leq m$. Also, $\operatorname{spo}_{\lambda}(X/\emptyset)=\mathrm{sp}_{\lambda}(X)$. Moreover,
-$\operatorname{spo}_{\lambda}(X/Y)=\sum_{\mu}\mathrm{sp}_{\mu}(X)s_{\lambda^{\prime}/\mu^{\prime}}(Y).$
+$\operatorname{spo}_{\lambda}(X;Y)\coloneqq\sum_{T\,\in\,\operatorname{OSP}_{n/m}({\lambda})}\operatorname{wt}(T).$
+Note that $\operatorname{spo}_{\lambda}(X;\emptyset)=\mathrm{sp}_{\lambda}(X)$. Moreover,
+$\operatorname{spo}_{\lambda}(X;Y)=\sum_{\mu}\mathrm{sp}_{\mu}(X)\mathrm{s}_{\lambda^{\prime}/\mu^{\prime}}(Y).$
 Example 2.7.
-Let $n=1$, $m=2$ and consider the partition $\lambda=(2)$. Then we have the following orthosymplectic tableaux in
+Let $n=1$, $m=2$ and consider the partition $\lambda=(2)$. Then we have the following orthosymplectic tableaux:
 […]
@@
 Therefore,
-$\operatorname{spo}_{(2)}(x_{1}/y_{1},y_{2})=x_{1}^{2}+\bar{x}_{1}^{2}+1+y_{1}(x_{1}+\bar{x}_{1})+y_{2}(x_{1}+\bar{x}_{1})+y_{1}y_{2}.$
+(2.12) $\operatorname{spo}_{(2)}(x_{1};y_{1},y_{2})=x_{1}^{2}+\bar{x}_{1}^{2}+1+y_{1}(x_{1}+\bar{x}_{1})+y_{2}(x_{1}+\bar{x}_{1})+y_{1}y_{2}.$
 The Jacobi–Trudi type identity for the orthosymplectic Schur functions is given by
-(2.13) $\operatorname{spo}_{\lambda}(X/Y)=\det\left(J_{\lambda_{i}-i+j}+J_{\lambda_{i}-i-j+2}\right)_{n\times n},$
-where $J_{r}=\sum_{l=0}^{r}h_{l}(X,\bar{X})e_{r-l}(Y)$.
-Note that the orthosymplectic Schur polynomial $\operatorname{spo}_{\lambda}(X/Y)$ is nonzero if and only if $\lambda_{n+1}\leq m$.
-The following remark gives an expression for $\operatorname{spo}_{\lambda}(X/y)$ if $\lambda_{n+1}\leq 1$.
+(2.13) $\operatorname{spo}_{\lambda}(X;Y)={\det\left(\begin{array}[]{c|c}\left(J_{\lambda_{i}-i+1}\right)_{1\leq i\leq n}&(J_{\lambda_{i}-i+j}+J_{\lambda_{i}-i-j+2})_{\begin{subarray}{c}1\leq i\leq n\\ 2\leq j\leq n\end{subarray}}\end{array}\right)},$
+where $J_{r}=\sum_{l=0}^{r}h_{l}(X,\overline{X})e_{r-l}(Y)$.
+Here $h_{l}(X,\overline{X})$ is the Laurent polynomial obtained by substituting $x_{n+i}=\bar{x}_{i},1\leq i\leq n$ in $h_{l}({x_{1},x_{2},\dots,x_{2n}})$.
+Note that the orthosymplectic Schur polynomial $\operatorname{spo}_{\lambda}(X;Y)$ is nonzero if and only if $\lambda_{n+1}\leq m$.
+The following remark gives an expression for $\operatorname{spo}_{\lambda}(X;y)$ if $\lambda_{n+1}\leq 1$.
 Remark 2.8.
-[…]
-$\operatorname{spo}_{\lambda}(X/y)=\frac{y^{\ell-n}}{D}\det\left(x_{i}^{\lambda_{j}+n-j+1}-\bar{x}_{i}^{\lambda_{j}+n-j+1}+y(x_{i}^{\lambda_{j}+n-j}-\bar{x}_{i}^{\lambda_{j}+n-j})\right),$
+For $m=1$, suppose $\lambda$ is a partition such that $\ell(\lambda)=\ell$ and $\lambda_{n+1}\leq 1$.
+If $\ell\leq n$, then $\operatorname{spo}_{\lambda}(X;y)$ is given by 1.4.
+If $\ell>n$, then $\lambda_{n+1}=\cdots=\lambda_{\ell}=1$.
+So, by 2.5, $T$ contains $1^{\prime}$ in all the boxes of $\lambda$ from the $(n+1)^{\text{th}}$ row to $\ell^{\text{th}}$ row, for all $T\in\operatorname{OSP}_{n/1}(\lambda)$.
+Therefore, using (2.10) and (2.11) we have
+$\operatorname{spo}_{\lambda}(X;y)=\frac{y^{\ell-n}}{D}\det\left(x_{i}^{\lambda_{j}+n-j+1}-\bar{x}_{i}^{\lambda_{j}+n-j+1}+y(x_{i}^{\lambda_{j}+n-j}-\bar{x}_{i}^{\lambda_{j}+n-j})\right),$
 where $D=\prod_{i=1}^{n}(x_{i}-\bar{x}_{i})\prod_{1\leq i<j\leq n}(x_{i}+\bar{x}_{i}-x_{j}-\bar{x}_{j})$.
@@
 Definition 2.9.
-Let $\lambda$ be a partition of length at most $n$. A $(n-1,1)$-symplectic tableau of shape $\lambda$ is a filling of the Young diagram of $\lambda$ with entries from
+Let $\lambda$ be a partition of length at most $n$. An $(n-1,1)$-symplectic tableau of shape $\lambda$ is a filling of the Young diagram of $\lambda$ with entries from
 $1<\overline{1}<2<\overline{2}<\cdots<n-1<\overline{n-1}<n$
 satisfying the following three conditions:
-  • the entries in each row are weakly increasing;
-  • the entries in each column are strictly increasing;
-  • the entries of the $i$’th row are greater than or equal to $i$
-We denote by $\operatorname{SPT}^{(n-1,1)}(\lambda)$ the set of $(n-1,1)$-symplectic tableaux of shape $\lambda$. Given a $(n-1,1)$-symplectic tableau $T$, we define
+  • the entries increase weakly across rows and strictly down columns,
+  • the entries of the $i^{\text{th}}$ row are greater than or equal to $i$.
+We denote by ${\operatorname{SP}^{(n-1,1)}(\lambda)}$ the set of $(n-1,1)$-symplectic tableaux of shape $\lambda$. Given an $(n-1,1)$-symplectic tableau $T$, we define
 $\operatorname{wt}(T)=x_{n}^{\alpha_{n}(T)}\prod_{i=1}^{n-1}x_{i}^{\alpha_{i}(T)-\alpha_{\overline{i}}(T)},$
-where $\alpha_{i}(T)$, $\alpha_{\overline{i}}(T)$ for $i\in[n-1]$, and $\alpha_{n}(T)$ are number of occurences of $i$, $\overline{i}$ and $n$ respectively.
+where $\alpha_{i}(T)$, $\alpha_{\overline{i}}(T)$ for $i\in[n-1]$, and $\alpha_{n}(T)$ are the number of occurences of $i$, $\overline{i}$ and $n$ respectively.
 The odd symplectic character indexed by a partition of length at most $n$ is given by
-$\mathrm{sp}^{(n-1,1)}(x_{1},\dots,x_{n-1}|x_{n})=\sum_{T\in\operatorname{SPT}^{(n-1,1)}(\lambda)}\operatorname{wt}(T).$
+$\mathrm{sp}^{(n-1,1)}_{\lambda}(x_{1},\dots,x_{n-1}|x_{n})=\sum_{T\in\operatorname{SP}^{(n-1,1)}(\lambda)}\operatorname{wt}(T).$
+Example 2.10.
+Suppose $n=2$ and $\lambda=(2,1)$. Then we have the folowing $(1,1)$-symplectic tableaux of shape $\lambda$:
+[…]
+Therefore,
+$\mathrm{sp}^{(1,1)}_{(2,1)}(x_{1}|x_{2})=x_{1}^{2}x_{2}+x_{2}+x_{1}x_{2}^{2}+\bar{x}_{1}^{2}x_{2}+\bar{x}_{1}x_{2}^{2}.$
 Okada [Oka20] proved the following determinantal formula for odd symplectic characters.
 Theorem 2.11 ([Oka20, Theorem 1.1]).
-For a partition of length at most $n$, let $A_{\lambda}=(a_{ij})$ be a $n\times n$ matrix such that
+For a partition of length at most $n$, let $A_{\lambda}=(a_{ij})$ be an $n\times n$ matrix such that
 $a_{ij}=\begin{cases}(x_{i}^{\lambda_{j}+n-j+1}-\bar{x}_{i}^{\lambda_{j}+n-j+1})-y^{-1}(x_{i}^{\lambda_{j}+n-j}-\bar{x}_{i}^{\lambda_{j}+n-j})&1\leq i\leq n-1,\\ y^{\lambda_{j}+n-j}&i=n.\end{cases}$
@@
 $\det A_{\emptyset}=\prod_{i=1}^{n-1}(x_{i}-x_{i}^{-1})\prod_{1\leq i<j\leq n}(x_{i}+\bar{x}_{i}-x_{j}-\bar{x}_{j}).$
-Comparing the determinantal formulas in 1.4 and 2.10, we get the following result.
+Comparing the determinantal formulas in 1.4 and 2.11, we get the following result.
 Corollary 2.12.
 Suppose $\lambda$ is a partition of length at most $n$. Then
-$\mathrm{sp}^{(n-1,1)}_{\lambda}(x_{1},\dots,x_{n-1}|x_{n})=\operatorname{spo}_{\lambda}(x_{1},\dots,x_{n-1},x_{n}^{-1}/-x_{n}^{-1}).$
+$\mathrm{sp}^{(n-1,1)}_{\lambda}(x_{1},\dots,x_{n-1}|x_{n})=\operatorname{spo}_{\lambda}(x_{1},\dots,x_{n-1},x_{n}^{-1};-x_{n}^{-1}).$
 3. Determinant formula for orthosymplectic Schur functions
-In this section, we give a proof of 1.2. We first prove an equivalent statement of 1.2 given in 3.3. Before that, let us recall the following lemmas.
+In this section, we give a proof of 1.2. We first prove 3.3, which is equivalent to 1.2. Before that, let us recall the following lemmas.
 Lemma 3.1 ([FH91, Lemma A.54]).
 Let $\zeta_{j}(p)=x_{j}^{p}-\bar{x}_{j}^{p}$, $1\leq j\leq n$. Then for any integer $l\geq 0$, $\zeta_{j}(l)$ is the product of the $1\times n$, $n\times n$ and $n\times 1$ matrices
-$\left(\begin{array}[]{cccc}\overline{H}_{l-n}&\overline{H}_{l-n+1}+\overline{H}_{l-n-1}&\dots&\overline{H}_{l-1}+\overline{H}_{l-2n+1}\end{array}\right)\left(\begin{array}[]{c}(-1)^{q-p}e_{q-p}\end{array}\right)\left(\begin{array}[]{c}\zeta_{j}(n)\\ \zeta_{j}(n-1)\\ \vdots\\ \zeta_{j}(1)\end{array}\right),$
-where $\overline{H}_{r}=h_{r}(X,\bar{X})$ for all $r\in\mathbb{Z}$.
+$\left(\begin{array}[]{c|c}\overline{H}_{l-n}&\left(\overline{H}_{l-n-1+j}+\overline{H}_{l-n+1-j}\right)_{2\leq j\leq n}\end{array}\right)\Big{(}(-1)^{v-u}e_{v-u}\Big{)}_{1\leq u,v\leq n}\Big{(}\zeta_{j}(n+1-u)\Big{)}_{1\leq u\leq n},$
+where $\overline{H}_{r}=h_{r}(X,\overline{X})$ for all $r\in\mathbb{Z}$.
 Lemma 3.2 ([Mac15, Chapter I.1, Equation (1.7)]).
-Let $\lambda$ be a partition such that $\ell(\lambda)\leq n_{1}$ and $\lambda_{1}\leq n_{2}$. Then the $n_{1}+n_{2}$ numbers
-$\lambda_{i}+n_{1}-i\quad(1\leq i\leq n_{1}),\qquad\qquad n_{1}-1+j-\lambda_{j}^{\prime}\quad(1\leq j\leq n_{2})$
-are a permutation of $\{0,1,\dots,n_{1}+n_{2}-1\}$.
+Let $\lambda$ be a partition such that $\ell(\lambda)\leq n_{1}$ and $\lambda_{1}\leq n_{2}$.
+Then $\{\lambda_{i}+n_{1}-i\mid 1\leq i\leq n_{1}\}$ and $\{n_{1}-1+j-\lambda_{j}^{\prime}\mid 1\leq j\leq n_{2}\}$ form a set partition of $\{0,1,\dots,n_{1}+n_{2}-1\}$.
 Theorem 3.3.
-Let $\lambda$ be a partition of length at most $n$ and $k=\min\{j|\lambda_{j}+n+1-j\leq m\}.$ Then the orthosymplectic Schur function $\operatorname{spo}_{\lambda}(X/Y)$ is given by
-(3.1) $\operatorname{spo}_{\lambda}(X/Y)=\frac{(-1)^{mn-n+k-1}}{D}\det\left(\begin{array}[]{c|c}R&X_{\lambda}\\[-8.5359pt] \\ \hline\cr\\[-8.5359pt] Y_{\lambda}&0\end{array}\right),$
+Let $\lambda$ be a partition of length at most $n$ and $k=\min\{j\mid\lambda_{j}+n+1-j\leq m\}.$ Then the orthosymplectic Schur function $\operatorname{spo}_{\lambda}(X;Y)$ is given by
+(3.1) $\operatorname{spo}_{\lambda}(X;Y)=\frac{(-1)^{mn-n+k-1}}{D}\det\left(\begin{array}[]{c|c}R&X_{\lambda}\\[-8.5359pt] \\ \hline\cr\\[-8.5359pt] Y_{\lambda}&0\end{array}\right),$
 where
@@
 $X_{\lambda}=\left({x_{i}^{\lambda_{j}+n-m-j+1}}{\prod_{q=1}^{m}(x_{i}+y_{q})}-{\bar{x}_{i}^{\lambda_{j}+n-m-j+1}}{\prod_{q=1}^{m}(\bar{x}_{i}+y_{q})}\right)_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq k-1\end{subarray}},$
-and
+$\psi$ and
 $D={\prod_{i=1}^{n}(x_{i}-\bar{x}_{i})\prod_{1\leq i<j\leq n}(x_{i}+\bar{x}_{i}-x_{j}-\bar{x}_{j})}.$
 Proof.
-By Jacobi–Trudi type identity for orthosymplectic Schur functions (2.13), we have
-(3.2) $\operatorname{spo}_{\lambda}(X/Y)=\det\left(J_{\lambda_{i}-i+j}+J_{\lambda_{i}-i-j+2}\right)_{1\leq i,j\leq n},$
-where $J_{r}=\displaystyle\sum_{p=0}^{r}h_{p}(X,\bar{X})e_{r-p}(Y)=\displaystyle\sum_{p=0}^{r}h_{r-p}(X,\bar{X})e_{p}(Y)$ for all $r\geq 0$.
-Otherwise $J_{r}=0$.
-Since $e_{p}(Y)=0$ for $p>m$ and $h_{r-p}(X,\bar{X})=0$, for $p>r$,
-(3.3) $J_{r}=\displaystyle\sum_{p=0}^{m}h_{r-p}(X,\bar{X})e_{p}(Y).,\quad r\in\mathbb{Z}.$
-Substituting the expression for $J_{r}$ from (3.3) in (3.2), we have
-$\operatorname{spo}_{\lambda}(X/Y)=\det\left(\sum_{p=0}^{m}(\overline{H}_{\lambda_{i}-i+j-p}+\overline{H}_{\lambda_{i}-i-j+2-p})e_{p}(Y)\right)_{n\times n},$
-where $\overline{H}_{r}=h_{r}(X,\bar{X})$ for all $r\in\mathbb{Z}$. Let $k=\text{min}\{j|\lambda_{j}+n+1-j\leq m\}.$ Now applying 3.1, the required orthosymplectic Schur function is given by
-$\frac{1}{D}\det\left(\begin{array}[]{c}\left(\displaystyle\sum_{p=0}^{m}\left(x_{j}^{\lambda_{i}+n-i+1-p}-\bar{x}_{j}^{\lambda_{i}+n-i+1-p}\right)e_{p}(Y)\right)_{1\leq i\leq k-1}\\[-8.5359pt] \\ \hline\cr\\[-8.5359pt] \left(\displaystyle\sum_{p=0}^{\lambda_{i}+n-i}\left(x_{j}^{\lambda_{i}+n-i+1-p}-\bar{x}_{j}^{\lambda_{i}+n-i+1-p}\right)e_{p}(Y)\right)_{k\leq i\leq n}\end{array}\right)_{1\leq j\leq n},$
+By the Jacobi–Trudi type identity for orthosymplectic Schur functions (2.13), we have
+(3.2) $\operatorname{spo}_{\lambda}(X;Y)={\det\left(\begin{array}[]{c|c}\left(J_{\lambda_{i}-i+1}\right)_{1\leq i\leq n}&(J_{\lambda_{i}-i+j}+J_{\lambda_{i}-i-j+2})_{\begin{subarray}{c}1\leq i\leq n\\ 2\leq j\leq n\end{subarray}}\end{array}\right)},$
+where $J_{r}=\displaystyle\sum_{p=0}^{r}h_{p}(X,\overline{X})e_{r-p}(Y)=\displaystyle\sum_{p=0}^{r}h_{r-p}(X,\overline{X})e_{p}(Y)$ for all $r\geq 0$ and $J_{r}=0$ otherwise.
+Since $e_{p}(Y)=0$ for $p>m$ and $h_{r-p}(X,\overline{X})=0$, for $p>r$,
+(3.3) $J_{r}=\displaystyle\sum_{p=0}^{m}h_{r-p}(X,\overline{X})e_{p}(Y),\quad r\in\mathbb{Z}.$
+Substituting the expression for $J_{r}$ from (3.3) in (3.2), the required orthosymplectic Schur function is given by
+$\det\left(\begin{array}[]{c|c}\left(\displaystyle\sum_{p=0}^{m}\overline{H}_{\lambda_{i}-i+1-p}e_{p}(Y)\right)_{1\leq i\leq n}&\left(\displaystyle\sum_{p=0}^{m}(\overline{H}_{\lambda_{i}-i+j-p}+\overline{H}_{\lambda_{i}-i-j+2-p})e_{p}(Y)\right)_{\begin{subarray}{c}1\leq i\leq n\\ 2\leq j\leq n\end{subarray}}\end{array}\right)$
+where $\overline{H}_{r}=h_{r}(X,\overline{X})$ for all $r\in\mathbb{Z}$. Observe that
+(3.4) $\operatorname{spo}_{\lambda}(X;Y)=\det\left(\sum_{p=0}^{m}e_{p}(Y)M^{p}\right),$
 where
-$D=\displaystyle\prod_{i=1}^{n}(x_{i}-\bar{x}_{i})\prod_{1\leq i<j\leq n}(x_{i}+\bar{x}_{i}-x_{j}-\bar{x}_{j}).$
+$M^{p}=\left(\begin{array}[]{c|c}\left(\displaystyle\overline{H}_{\lambda_{i}-i+1-p}\right)_{1\leq i\leq n}&\left(\displaystyle\overline{H}_{\lambda_{i}-i+j-p}+\overline{H}_{\lambda_{i}-i-j+2-p}\right)_{\begin{subarray}{c}1\leq i\leq n\\ 2\leq j\leq n\end{subarray}}\end{array}\right).$
+If $\lambda_{i}+n-i+1-p\geq 0$, then by 3.1, we have
+$(M^{p}_{i,v})_{1\leq v\leq n}\left((-1)^{v-u}e_{v-u}\right)_{1\leq u,v\leq n}\left(x_{j}^{n+1-u}-\bar{x}_{j}^{n+1-u}\right)_{1\leq u\leq n}=x_{j}^{\lambda_{i}+n-i+1-p}-\bar{x}_{j}^{\lambda_{i}+n-i+1-p}$
+and $M^{p}_{i,v}=0$ for all $1\leq v\leq n$ otherwise. As a consequence, we have
+(3.5) $\begin{split}\left(\sum_{p=0}^{m}e_{p}(Y)M^{p}\right)&\Big{(}(-1)^{v-u}e_{v-u}\Big{)}_{1\leq u,v\leq n}\Big{(}x_{j}^{n+1-u}-\bar{x}_{j}^{n+1-u}\Big{)}_{1\leq u\leq n}\\ =&\left(\begin{array}[]{c}\left(\displaystyle\sum_{p=0}^{m}\left(x_{j}^{\lambda_{i}+n-i+1-p}-\bar{x}_{j}^{\lambda_{i}+n-i+1-p}\right)e_{p}(Y)\right)_{1\leq i\leq k-1}\\[-8.5359pt] \\ \hline\cr\\[-8.5359pt] \left(\displaystyle\sum_{p=0}^{\lambda_{i}+n-i}\left(x_{j}^{\lambda_{i}+n-i+1-p}-\bar{x}_{j}^{\lambda_{i}+n-i+1-p}\right)e_{p}(Y)\right)_{k\leq i\leq n}\end{array}\right)_{n\times 1}.\end{split}$
+Therefore, using the expressions given in (3.4),(3.5) and the fact that
+$\det.\left((-1)^{v-u}e_{v-u}\right)_{1\leq u,v\leq n}=1,$
+the orthosymplectic Schur function $\operatorname{spo}_{\lambda}(X;Y)$ is given by
+(3.6) $\frac{1}{D}\det\left(\begin{array}[]{c}\left(\displaystyle\sum_{p=0}^{m}\left(x_{j}^{\lambda_{i}+n-i+1-p}-\bar{x}_{j}^{\lambda_{i}+n-i+1-p}\right)e_{p}(Y)\right)_{1\leq i\leq k-1}\\[-8.5359pt] \\ \hline\cr\\[-8.5359pt] \left(\displaystyle\sum_{p=0}^{\lambda_{i}+n-i}\left(x_{j}^{\lambda_{i}+n-i+1-p}-\bar{x}_{j}^{\lambda_{i}+n-i+1-p}\right)e_{p}(Y)\right)_{k\leq i\leq n}\end{array}\right)_{1\leq j\leq n},$
+where
+$D=\det\Big{(}x_{j}^{n+1-u}-\bar{x}_{j}^{n+1-u}\Big{)}_{1\leq u,j\leq n}=\displaystyle\prod_{i=1}^{n}(x_{i}-\bar{x}_{i})\prod_{1\leq i<j\leq n}(x_{i}+\bar{x}_{i}-x_{j}-\bar{x}_{j}).$
 Note that
 (3.7) $\sum_{p=0}^{m}x_{i}^{-p}e_{p}(Y)=\prod_{j=1}^{m}(1+\bar{x}_{i}y_{j}),\qquad\sum_{p=0}^{m}x_{i}^{p}e_{p}(Y)=\prod_{i=1}^{m}(1+x_{i}y_{j}).$
-Consider the permutation $\sigma=(k,\dots,n,1,\dots,k-1)$, in one line notation in $S_{n}$. Then, the number of inversions of $\sigma$ is $(k-1)(n-k+1)$ and it can be seen that
-(3.5) $\operatorname{spo}_{\lambda}(X/Y)=\frac{\operatorname{sgn}(\sigma)}{D}\det\left(\begin{array}[]{c|c}(A_{i,j})_{\begin{subarray}{c}1\leq i\leq n\\ k\leq j\leq n\end{subarray}}&(B_{i,j})_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq k-1\end{subarray}}\end{array}\right),$
+Consider the permutation $\sigma=(k,\dots,n,1,\dots,k-1)\in$ $S_{n}$.
+The number of inversions of $\sigma$ is $(k-1)(n-k+1)$.
+Permuting the rows of the determinant in (3.6) by $\sigma$ and then transposing the determinant, we have
+(3.8) $\operatorname{spo}_{\lambda}(X;Y)=\frac{\operatorname{sgn}(\sigma)}{D}\det\left(\begin{array}[]{c|c}(A_{i,j})_{\begin{subarray}{c}1\leq i\leq n\\ k\leq j\leq n\end{subarray}}&(B_{i,j})_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq k-1\end{subarray}}\end{array}\right),$
 where
@@
 where the last equality uses the relations given in (3.7).
-Let $\Lambda\coloneqq\{\lambda_{j}+n-j+1:k\leq j\leq n\}$ and $\Lambda^{\prime}\coloneqq\{n+i-\lambda_{i}^{\prime}:1\leq i\leq m-n+k-1\}$.
-Note that $\{\Lambda,\Lambda^{\prime}\}$ is a set partition of $\{1,2,\dots,m\}$ by 3.2.
-So, we can rewrite (3.5) in the following manner.
-$\operatorname{spo}_{\lambda}(X/Y)=\frac{(-1)^{b}}{D}\det\left(\begin{array}[]{c|c|c}((-1)^{m-j}(x_{i}^{j}-\bar{x}_{i}^{j}))_{\begin{subarray}{c}1\leq i\leq n\\ j\,\in\,\Lambda^{\prime}\end{subarray}}&(A_{i,j})_{\begin{subarray}{c}1\leq i\leq n\\ k\leq j\leq n\end{subarray}}&(B_{i,j})_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq k-1\end{subarray}}\\ &&\\ \hline\cr&\\ (h_{-i+j}(Y))_{\begin{subarray}{c}i\,\in\,\Lambda^{\prime}\\ j\,\in\,\Lambda^{\prime}\end{subarray}}&{\text{\huge 0}}&{\text{\huge 0}}\end{array}\right),$
+Let $\Lambda\coloneqq\{\lambda_{j}+n-j+1\mid k\leq j\leq n\}$, $\Lambda^{\prime}\coloneqq\{n+i-\lambda_{i}^{\prime}\mid 1\leq i\leq m-n+k-1\}$ and
+$M\coloneqq\left(\begin{array}[]{c|c|c}((-1)^{m-j}(x_{i}^{j}-\bar{x}_{i}^{j}))_{\begin{subarray}{c}1\leq i\leq n\\ j\,\in\,\Lambda^{\prime}\end{subarray}}&(A_{i,j})_{\begin{subarray}{c}1\leq i\leq n\\ k\leq j\leq n\end{subarray}}&(B_{i,j})_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq k-1\end{subarray}}\\ &&\\ \hline\cr&\\ (h_{-i+j}(Y))_{\begin{subarray}{c}i\,\in\,\Lambda^{\prime}\\ j\,\in\,\Lambda^{\prime}\end{subarray}}&{\text{\huge 0}}&{\text{\huge 0}}\end{array}\right).$
+Since
+$\det(h_{-i+j}(Y))_{\begin{subarray}{c}i\,\in\,\Lambda^{\prime}\\ j\,\in\,\Lambda^{\prime}\end{subarray}}=1,$
+$\det M=(-1)^{n(m-n+k-1)}\det\left(\begin{array}[]{c|c}(A_{i,j})_{\begin{subarray}{c}1\leq i\leq n\\ k\leq j\leq n\end{subarray}}&(B_{i,j})_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq k-1\end{subarray}}\end{array}\right).$
+So, we can rewrite (3.8) in the following manner:
+$\operatorname{spo}_{\lambda}(X;Y)=\frac{(-1)^{b}}{D}\det\left(\begin{array}[]{c|c|c}((-1)^{m-j}(x_{i}^{j}-\bar{x}_{i}^{j}))_{\begin{subarray}{c}1\leq i\leq n\\ j\,\in\,\Lambda^{\prime}\end{subarray}}&(A_{i,j})_{\begin{subarray}{c}1\leq i\leq n\\ k\leq j\leq n\end{subarray}}&(B_{i,j})_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq k-1\end{subarray}}\\ &&\\ \hline\cr&\\ (h_{-i+j}(Y))_{\begin{subarray}{c}i\,\in\,\Lambda^{\prime}\\ j\,\in\,\Lambda^{\prime}\end{subarray}}&{\text{\huge 0}}&{\text{\huge 0}}\end{array}\right),$
 where $b=(k-1)(n-k+1)+n(m-n+k-1)=mn-(n-k+1)^{2}$. Observe that
 $(A_{i,j})_{\begin{subarray}{c}1\leq i\leq n\\ k\leq j\leq n\end{subarray}}=\left(\sum_{p=0}^{j-1}(x_{i}^{j-p}-\bar{x}_{i}^{j-p})e_{p}(Y)\right)_{\begin{subarray}{c}1\leq i\leq n\\ j\,\in\,\Lambda\end{subarray}}$
-So, the orthosymplectic Schur function $\operatorname{spo}_{\lambda}(X/Y)$ is given by
+So, the orthosymplectic Schur function $\operatorname{spo}_{\lambda}(X;Y)$ is given by
 (3.9) $\frac{(-1)^{b}}{D}\det\left(\begin{array}[]{c|c|c}\left((-1)^{m-j}\phi_{i}(j)\right)_{\begin{subarray}{c}1\leq i\leq n\\ j\,\in\,\Lambda^{\prime}\end{subarray}}&\left(\sum_{p=0}^{j-1}\phi_{i}(j-p)e_{p}(Y)\right)_{\begin{subarray}{c}1\leq i\leq n\\ j\,\in\,\Lambda\end{subarray}}&(B_{i,j})_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq k-1\end{subarray}}\\ &&\\ \hline\cr&\\ (h_{-i+j}(Y))_{\begin{subarray}{c}i\,\in\,\Lambda^{\prime}\\ j\,\in\,\Lambda^{\prime}\end{subarray}}&{\text{\huge 0}}&{\text{\huge 0}}\end{array}\right),$
-where $\phi_{i}(j)=x_{i}^{j}-\bar{x}_{i}^{j}$. Applying $C_{\lambda_{n}+1}\rightarrow C_{\lambda_{n}+1}-\sum_{p=1}^{\lambda_{n}}(-1)^{m-p}e_{\lambda_{n}+1-p}(Y)C_{p}$ and using
+where $\phi_{i}(j)=x_{i}^{j}-\bar{x}_{i}^{j}$.
+Note that $\{\Lambda,\Lambda^{\prime}\}$ is a set partition of $\{1,2,\dots,m\}$ by 3.2.
+Applying the elementary column transformation $C_{\lambda_{n}+1}\rightarrow C_{\lambda_{n}+1}-\sum_{p=1}^{\lambda_{n}}(-1)^{m-p}e_{\lambda_{n}+1-p}(Y)C_{p}$ and using
 $-\sum_{p=1}^{\lambda_{n}}(-1)^{m-p}e_{\lambda_{n}+1-p}(Y)h_{-i+p}(Y)=(-1)^{m-\lambda_{n}+1}h_{-i+\lambda_{n}+1}(Y),$
-the function in (3.6) is the same as
-$\frac{(-1)^{b}}{D}\\ \times\left|\begin{array}[]{c|c|c|c}\left((-1)^{m-j}\phi_{i}(j)\right)_{\begin{subarray}{c}1\leq i\leq n\\ j\,\in\,\Lambda^{\prime}\end{subarray}}&\left(\displaystyle\sum_{p=0}^{j-1}\psi_{i}(p)e_{p}(Y)\right)_{\begin{subarray}{c}1\leq i\leq n\\ j\,\in\,\Lambda\end{subarray}}&(\phi_{i}(j))_{\begin{subarray}{c}1\leq i\leq n\\ j=\lambda_{n}+1\end{subarray}}&(B_{i,j})_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq k-1\end{subarray}}\\ &&&\\ \hline\cr&&\\ (h_{-i+j}(Y))_{\begin{subarray}{c}i\,\in\,\Lambda^{\prime}\\ j\,\in\,\Lambda^{\prime}\end{subarray}}&{\text{\huge 0}}&((-1)^{m-j}h_{-i+j}(Y))_{\begin{subarray}{c}i\in\Lambda^{\prime}\\ j=\lambda_{n}+1\end{subarray}}&{\text{\huge 0}}\end{array}\right|,$
-where $\psi_{i}(p)=\phi_{i}(j-p)$. Successively applying $C_{j}\rightarrow C_{j}-\sum_{p=1}^{j-1}(-1)^{m-p}e_{j-p}C_{p}$, for $j\in\Lambda$, we have
-$\operatorname{spo}_{\lambda}(X/Y)=\frac{(-1)^{b}}{D}\det\left(\begin{array}[]{c|c|c}\left((-1)^{m-j}\phi(j)\right)_{\begin{subarray}{c}1\leq i\leq n\\ j\,\in\,\Lambda^{\prime}\end{subarray}}&(\phi(j))_{\begin{subarray}{c}1\leq i\leq n\\ j\,\in\,\Lambda\end{subarray}}&(B_{i,j})_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq k-1\end{subarray}}\\ &&\\ \hline\cr&\\ (h_{-i+j}(Y))_{\begin{subarray}{c}i\,\in\,\Lambda^{\prime}\\ j\,\in\,\Lambda^{\prime}\end{subarray}}&((-1)^{m-j}h_{-i+j}(Y))_{\begin{subarray}{c}i\,\in\,\Lambda^{\prime}\\ j\,\in\,\Lambda\end{subarray}}&{\text{\huge 0}}\end{array}\right).$
-Rearranging the columns, we have
-$\operatorname{spo}_{\lambda}(X/Y)=\frac{(-1)^{a}}{D}\det\left(\begin{array}[]{c|c}\left((-1)^{m-j}(x_{i}^{j}-\bar{x}_{i}^{j})\right)_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq m\end{subarray}}&(B_{i,j})_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq k-1\end{subarray}}\\[17.50002pt] \hline\cr\\ \left(h_{\lambda^{\prime}_{i}-n-i+j}(Y)\right)_{\begin{subarray}{c}1\leq i\leq m-n+k-1\\ 1\leq j\leq m\end{subarray}}&{\text{\huge 0}}\end{array}\right),$
+the expression in (3.9) is the same as
+${\footnotesize\frac{(-1)^{b}}{D}\left|\begin{array}[]{c|c|c|c}\left((-1)^{m-j}\phi_{i}(j)\right)_{\begin{subarray}{c}1\leq i\leq n\\ j\,\in\,\Lambda^{\prime}\end{subarray}}&\left(\displaystyle\sum_{p=0}^{j-1}\phi_{i}(j-p)e_{p}(Y)\right)_{\begin{subarray}{c}1\leq i\leq n\\ j\,\in\,\Lambda\\ j\neq\lambda_{n}+1\end{subarray}}&(\phi_{i}(j))_{\begin{subarray}{c}1\leq i\leq n\\ j=\lambda_{n}+1\end{subarray}}&(B_{i,j})_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq k-1\end{subarray}}\\ &&&\\ \hline\cr&&\\ (h_{-i+j}(Y))_{\begin{subarray}{c}i\,\in\,\Lambda^{\prime}\\ j\,\in\,\Lambda^{\prime}\end{subarray}}&{\text{\huge 0}}&\left((-1)^{m-j}h_{-i+j}(Y)\right)_{\begin{subarray}{c}i\in\Lambda^{\prime}\\ j=\lambda_{n}+1\end{subarray}}&{\text{\huge 0}}\end{array}\right|.}$
+Successively applying the elementary column transformations $C_{j}\rightarrow C_{j}-\sum_{p=1}^{j-1}(-1)^{m-p}e_{j-p}C_{p}$, for $j\in\Lambda$, we have
+$\operatorname{spo}_{\lambda}(X;Y)=\frac{(-1)^{b}}{D}\det\left(\begin{array}[]{c|c|c}\left((-1)^{m-j}\phi(j)\right)_{\begin{subarray}{c}1\leq i\leq n\\ j\,\in\,\Lambda^{\prime}\end{subarray}}&(\phi(j))_{\begin{subarray}{c}1\leq i\leq n\\ j\,\in\,\Lambda\end{subarray}}&(B_{i,j})_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq k-1\end{subarray}}\\ &&\\ \hline\cr&\\ (h_{-i+j}(Y))_{\begin{subarray}{c}i\,\in\,\Lambda^{\prime}\\ j\,\in\,\Lambda^{\prime}\end{subarray}}&((-1)^{m-j}h_{-i+j}(Y))_{\begin{subarray}{c}i\,\in\,\Lambda^{\prime}\\ j\,\in\,\Lambda\end{subarray}}&{\text{\huge 0}}\end{array}\right).$
+Taking $(-1)^{m-j}$ out from the $j^{\text{th}}$ column for each $j\in\Lambda$, the orthosymplectic Schur function is given by
+$\operatorname{spo}_{\lambda}(X;Y)=\frac{(-1)^{b^{\prime}}}{D}\det\left(\begin{array}[]{c|c|c}\left((-1)^{m-j}\phi(j)\right)_{\begin{subarray}{c}1\leq i\leq n\\ j\,\in\,\Lambda^{\prime}\end{subarray}}&((-1)^{m-j}\phi(j))_{\begin{subarray}{c}1\leq i\leq n\\ j\,\in\,\Lambda\end{subarray}}&(B_{i,j})_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq k-1\end{subarray}}\\ &&\\ \hline\cr&\\ (h_{-i+j}(Y))_{\begin{subarray}{c}i\,\in\,\Lambda^{\prime}\\ j\,\in\,\Lambda^{\prime}\end{subarray}}&(h_{-i+j}(Y))_{\begin{subarray}{c}i\,\in\,\Lambda^{\prime}\\ j\,\in\,\Lambda\end{subarray}}&{\text{\huge 0}}\end{array}\right).$
+where $b^{\prime}=b+\sum_{j\in\Lambda}(m-j)$. Rearranging the columns, we have
+$\operatorname{spo}_{\lambda}(X;Y)=\frac{(-1)^{a}}{D}\det\left(\begin{array}[]{c|c}\left((-1)^{m-j}(x_{i}^{j}-\bar{x}_{i}^{j})\right)_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq m\end{subarray}}&(B_{i,j})_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq k-1\end{subarray}}\\[17.50002pt] \hline\cr\\ \left(h_{\lambda^{\prime}_{i}-n-i+j}(Y)\right)_{\begin{subarray}{c}1\leq i\leq m-n+k-1\\ 1\leq j\leq m\end{subarray}}&{\text{\huge 0}}\end{array}\right),$
 where $a=mn-n+k-1$. This completes the proof. ∎
-Now we prove that the expressions for $\operatorname{spo}_{\lambda}(X/Y)$ in 1.2 and 3.3 are the same.
+Now we prove that the expressions for $\operatorname{spo}_{\lambda}(X;Y)$ in 1.2 and 3.3 are the same.
 Proof of 1.2.
-Assume that (1.1) holds true. Let $a=mn-n+k-1$. Then the orthosymplectic Schur function $\operatorname{spo}_{\lambda}(X/Y)$ is given by
-(3.7) $\frac{(-1)^{a}}{D_{s}}\det\left(\begin{array}[]{c|c}\left(\displaystyle\frac{x_{i}}{(x_{i}+y_{j})\prod_{q=1}^{m}(\bar{x}_{i}+y_{q})}-\displaystyle\frac{\bar{x}_{i}}{(\bar{x}_{i}+y_{j})\prod_{q=1}^{m}(x_{i}+y_{q})}\right)_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq m\end{subarray}}&X_{\lambda}\\[17.50002pt] \hline\cr\\ \left(y_{j}^{\lambda^{\prime}_{i}+m-n-i}\right)_{\begin{subarray}{c}1\leq i\leq m-n+k-1\\ 1\leq j\leq m\end{subarray}}&{\text{\huge 0}}\end{array}\right),$
+Let $a=mn-n+k-1$. Consider the following expression given in (1.1).
+${A}\coloneqq\frac{(-1)^{a}}{D_{s}}\det\left(\begin{array}[]{c|c}\left(\displaystyle\frac{x_{i}}{(x_{i}+y_{j})\prod_{q=1}^{m}(\bar{x}_{i}+y_{q})}-\displaystyle\frac{\bar{x}_{i}}{(\bar{x}_{i}+y_{j})\prod_{q=1}^{m}(x_{i}+y_{q})}\right)_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq m\end{subarray}}&X_{\lambda}\\[17.50002pt] \hline\cr\\ \left(y_{j}^{\lambda^{\prime}_{i}+m-n-i}\right)_{\begin{subarray}{c}1\leq i\leq m-n+k-1\\ 1\leq j\leq m\end{subarray}}&{\text{\huge 0}}\end{array}\right),$
 where
@@
 $D_{s}=\frac{\prod_{i=1}^{n}(x_{i}-\bar{x}_{i})\prod_{1\leq i<j\leq n}(x_{i}+\bar{x}_{i}-x_{j}-\bar{x}_{j})\prod_{1\leq i<j\leq m}(y_{i}-y_{j})}{\prod_{i=1}^{n}\prod_{j=1}^{m}(x_{i}+y_{j})(\bar{x}_{i}+y_{j})}.$
-Applying $R_{i}\rightarrow\prod_{q=1}^{m}(x_{i}+y_{q})(\bar{x}_{i}+y_{q})R_{i}$, for $i\in[1,n]$ and cancelling the term $\prod_{i=1}^{n}\prod_{j=1}^{m}(x_{i}+y_{j})(\bar{x}_{i}+y_{j})$ from the denominator of $D_{s}$, (3.7) can be written as
+Applying the elementary row transformations $R_{i}\rightarrow\prod_{q=1}^{m}(x_{i}+y_{q})(\bar{x}_{i}+y_{q})R_{i}$, for $i\in[1,n]$ and cancelling the term $\prod_{i=1}^{n}\prod_{j=1}^{m}(x_{i}+y_{j})(\bar{x}_{i}+y_{j})$ from the denominator of $D_{s}$, the expression for $A$ can be written as
 $\frac{(-1)^{a}}{D_{s}^{1}}\det\left(\begin{array}[]{c|c}\left(x_{i}\prod_{\begin{subarray}{c}q=1\\ q\neq j\end{subarray}}^{m}(x_{i}+y_{q})-\bar{x}_{i}\prod_{\begin{subarray}{c}q=1\\ q\neq j\end{subarray}}^{m}(\bar{x}_{i}+y_{q})\right)_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq m\end{subarray}}&Z_{\lambda}\\[17.50002pt] \hline\cr\\ \left(y_{j}^{\lambda^{\prime}_{i}+m-n-i}\right)_{\begin{subarray}{c}1\leq i\leq m-n+k-1\\ 1\leq j\leq m\end{subarray}}&{\text{\huge 0}}\end{array}\right),$
@@
 and
-$Z_{\lambda}=\left({x_{i}^{\lambda_{j}+n-m-j+1}}{\prod_{q=1}^{m}(x_{i}+y_{q})}-{\bar{x}_{i}^{\lambda_{j}+n-m-j+1}}{\prod_{q=1}^{m}(\bar{x}_{i}+y_{q})}\right)_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq k-1\end{subarray}},$
+$Z_{\lambda}=\left({x_{i}^{\lambda_{j}+n-m-j+1}}{\prod_{q=1}^{m}(x_{i}+y_{q})}-{\bar{x}_{i}^{\lambda_{j}+n-m-j+1}}{\prod_{q=1}^{m}(\bar{x}_{i}+y_{q})}\right)_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq k-1\end{subarray}}.$
 Let
 $M^{k}_{i,j}=x_{i}\prod_{\begin{subarray}{c}q=1\\ q\neq j\end{subarray}}^{k}(x_{i}+y_{q})-\bar{x}_{i}\prod_{\begin{subarray}{c}q=1\\ q\neq j\end{subarray}}^{k}(\bar{x}_{i}+y_{q}).$
-Applying $C_{j}\rightarrow C_{j}-C_{m}$, $j\in[1,m-1]$ and then using $y_{j}^{p}-y_{m}^{p}=(y_{j}-y_{m})h_{p-1}(y_{j},y_{m})$, the orthosymplectic Schur function $\operatorname{spo}_{\lambda}(X/Y)$ is given by
+Applying the elementary column transformations $C_{j}\rightarrow C_{j}-C_{m}$, $j\in[1,m-1]$ and then using $y_{j}^{p}-y_{m}^{p}=(y_{j}-y_{m})h_{p-1}(y_{j},y_{m})$ to factor out the term $y_{j}-y_{m}$ for $j\in[1,m-1]$, $A$ is given by
 $\frac{(-1)^{a}}{D_{s}^{2}}\det\left(\begin{array}[]{c|c|c}\left(-M^{m-1}_{i,j}\right)_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq m-1\end{subarray}}&\left(M^{m}_{i,j}\right)_{1\leq i\leq n}&Z_{\lambda}\\ &\\ \hline\cr&\\ \left(h_{\lambda^{\prime}_{i}+m-n-i-1}(y_{j},y_{m})\right)_{\begin{subarray}{c}1\leq i\leq m-n+k-1\\ 1\leq j\leq m-1\end{subarray}}&\left(h_{\lambda^{\prime}_{i}+m-n-i}(y_{m})\right)_{1\leq i\leq m-n+k-1}&{\text{\huge 0}}\end{array}\right),$
@@
 $D_{s}^{2}=\prod_{i=1}^{n}(x_{i}-\bar{x}_{i})\prod_{1\leq i<j\leq n}(x_{i}+\bar{x}_{i}-x_{j}-\bar{x}_{j})\prod_{1\leq i<j\leq m-1}(y_{i}-y_{j}).$
-Now applying $C_{j}\rightarrow C_{j}-C_{s}$, for $j\in[1,s-1]$ and then using $h_{p}(y_{j},y_{s+1},y_{s+2},\dots,y_{m})-h_{p}(y_{s},y_{s+1},y_{s+2},\dots,y_{m})=(y_{j}-y_{s})h_{p-1}(y_{j},y_{s},y_{s+1},y_{s+2},\dots,y_{m})$, successively for $s=m-1,m-2,\dots,2$, we have
-$\operatorname{spo}_{\lambda}(X/Y)=\frac{(-1)^{a}}{D}\det\left(\begin{array}[]{c|c}\left((-1)^{m-j}(x_{i}\prod_{q=1}^{j-1}(x_{i}+y_{q})-\bar{x}_{i}\prod_{q=1}^{j-1}(\bar{x}_{i}+y_{q}))\right)_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq m\end{subarray}}&Z_{\lambda}\\[17.50002pt] \hline\cr\\ \left(h_{\lambda^{\prime}_{i}-n-i+j}(y_{j},\dots,y_{m})\right)_{\begin{subarray}{c}1\leq i\leq m-n+k-1\\ 1\leq j\leq m\end{subarray}}&{\text{\huge 0}}\end{array}\right),$
+Now applying the elementary column transformations $C_{j}\rightarrow C_{j}-C_{s}$, $j\in[1,s-1]$ and using $h_{p}(y_{j},y_{s+1},y_{s+2},\dots,y_{m})-h_{p}(y_{s},y_{s+1},y_{s+2},\dots,y_{m})=(y_{j}-y_{s})h_{p-1}(y_{j},y_{s},y_{s+1},y_{s+2},\dots,y_{m}),$ successively for $s=m-1,m-2,\dots,2$, we have
+$A=\frac{(-1)^{a}}{D}\det\left(\begin{array}[]{c|c}\left((-1)^{m-j}(x_{i}\prod_{q=1}^{j-1}(x_{i}+y_{q})-\bar{x}_{i}\prod_{q=1}^{j-1}(\bar{x}_{i}+y_{q}))\right)_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq m\end{subarray}}&Z_{\lambda}\\[17.50002pt] \hline\cr\\ \left(h_{\lambda^{\prime}_{i}-n-i+j}(y_{j},\dots,y_{m})\right)_{\begin{subarray}{c}1\leq i\leq m-n+k-1\\ 1\leq j\leq m\end{subarray}}&{\text{\huge 0}}\end{array}\right),$
 where
 $D=\prod_{i=1}^{n}(x_{i}-\bar{x}_{i})\prod_{1\leq i<j\leq n}(x_{i}+\bar{x}_{i}-x_{j}-\bar{x}_{j}).$
-Applying $C_{j}\rightarrow C_{j}+y_{j-1}C_{j-1}$, for $j\in[2,m]$ and using $h_{n}(y_{j-1},\dots,y_{m})=h_{n}(y_{j},\dots,y_{m})+y_{j-1}h_{n-1}(y_{j-1},\dots,y_{m})$, $\operatorname{spo}_{\lambda}(X/Y)$ is given by
+Applying the elementary column transformations $C_{j}\rightarrow C_{j}+y_{j-1}C_{j-1}$, for $j\in[2,m]$ and using $h_{n}(y_{j-1},\dots,y_{m})=h_{n}(y_{j},\dots,y_{m})+y_{j-1}h_{n-1}(y_{j-1},\dots,y_{m})$, $A$ is given by
 $\frac{(-1)^{a}}{D}\det\left(\begin{array}[]{c|c|c}\left((-1)^{m-1}(x_{i}-\bar{x}_{i})\right)_{1\leq i\leq n}&\left(N_{i,j}^{2}\right)_{\begin{subarray}{c}1\leq i\leq n\\ 2\leq j\leq m\end{subarray}}&Z_{\lambda}\\ &&\\ \hline\cr&&\\ \left(h_{\lambda^{\prime}_{i}-n-i+1}(Y)\right)_{1\leq i\leq m-n+k-1}&\left(h_{\lambda^{\prime}_{i}-n-i+j}(y_{j-1},\dots,y_{m})\right)_{\begin{subarray}{c}1\leq i\leq m-n+k-1\\ 2\leq j\leq m\end{subarray}}&{\text{\huge 0}}\end{array}\right),$
@@
 $N_{i,j}^{s}=(-1)^{m-j}\left(x_{i}^{s}\prod_{q=1}^{j-s}(x_{i}+y_{q})-\bar{x}_{i}^{s}\prod_{q=1}^{j-s}(\bar{x}_{i}+y_{q})\right).$
-Now again applying $C_{j}\rightarrow C_{j}+y_{j-1}C_{j-1}$, for $j\in[3,m]$, we see that the orthosymplectic Schur function is given by
-$\frac{(-1)^{a}}{D}\det\left(\begin{array}[]{c|c|c}\left((-1)^{m-j}(x_{i}^{j}-\bar{x}_{i}^{j})\right)_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq 2\end{subarray}}&\left(N_{i,j}^{3}\right)_{\begin{subarray}{c}1\leq i\leq n\\ 3\leq j\leq m\end{subarray}}&Z_{\lambda}\\ &&\\ \hline\cr&&\\ \left(h_{\lambda^{\prime}_{i}-n-i+j}(Y)\right)_{\begin{subarray}{c}1\leq i\leq m-n+k-1\\ 1\leq j\leq 2\end{subarray}}&\left(h_{\lambda^{\prime}_{i}-n-i+j}(y_{j-2},\dots,y_{m})\right)_{\begin{subarray}{c}1\leq i\leq m-n+k-1\\ 2\leq j\leq m\end{subarray}}&{\text{\huge 0}}\end{array}\right).$
+Now again applying the elementary column transformations $C_{j}\rightarrow C_{j}+y_{j-2}C_{j-1}$, for $j\in[3,m]$, we see that
+$A=\frac{(-1)^{a}}{D}\det\left(\begin{array}[]{c|c|c}\left((-1)^{m-j}(x_{i}^{j}-\bar{x}_{i}^{j})\right)_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq 2\end{subarray}}&\left(N_{i,j}^{3}\right)_{\begin{subarray}{c}1\leq i\leq n\\ 3\leq j\leq m\end{subarray}}&Z_{\lambda}\\ &&\\ \hline\cr&&\\ \left(h_{\lambda^{\prime}_{i}-n-i+j}(Y)\right)_{\begin{subarray}{c}1\leq i\leq m-n+k-1\\ 1\leq j\leq 2\end{subarray}}&\left(h_{\lambda^{\prime}_{i}-n-i+j}(y_{j-2},\dots,y_{m})\right)_{\begin{subarray}{c}1\leq i\leq m-n+k-1\\ 2\leq j\leq m\end{subarray}}&{\text{\huge 0}}\end{array}\right).$
 Proceeding in a similar fashion, we have
-(3.8) $\operatorname{spo}_{\lambda}(X/Y)=\frac{(-1)^{a}}{D}\det\left(\begin{array}[]{c|c}\left((-1)^{m-j}(x_{i}^{j}-\bar{x}_{i}^{j})\right)_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq m\end{subarray}}&Z_{\lambda}\\[17.50002pt] \hline\cr\\ \left(h_{\lambda^{\prime}_{i}-n-i+j}(Y)\right)_{\begin{subarray}{c}1\leq i\leq m-n+k-1\\ 1\leq j\leq m\end{subarray}}&{\text{\huge 0}}\end{array}\right).$
-Observe that the expression in (3.8) is the same as the expression in (3.1). This completes the proof. ∎
+(3.10) $A=\frac{(-1)^{a}}{D}\det\left(\begin{array}[]{c|c}\left((-1)^{m-j}(x_{i}^{j}-\bar{x}_{i}^{j})\right)_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq m\end{subarray}}&Z_{\lambda}\\[17.50002pt] \hline\cr\\ \left(h_{\lambda^{\prime}_{i}-n-i+j}(Y)\right)_{\begin{subarray}{c}1\leq i\leq m-n+k-1\\ 1\leq j\leq m\end{subarray}}&{\text{\huge 0}}\end{array}\right).$
+Observe that the expression in (3.10) is the same as the expression in (3.1), which is equal to $\operatorname{spo}_{\lambda}(X;Y)$. This completes the proof. ∎
 4. A generalization of Brent–Krattenthaler–Warnaar’s identity
 In this section, we prove a generalization of the following odd symplectic character identity due to Brent, Krattenthaler and Warnaar, which was found in the study of discrete Mehta-type integrals [BKW16].
-Recently, Okada gave its linear-algebraic proof [Oka20].
+Recently, Okada gave a linear-algebraic proof [Oka20].
 Theorem 4.1 ([Oka20, Theorem 1.2]).
@@
 $\sum_{\lambda}z^{-r}\mathrm{sp}^{(n,1)}_{\lambda}(X|z)\,\mathrm{sp}^{(m,1)}_{(r^{m-n})\cup\lambda}(Y|z)=\mathrm{sp}_{(\underbrace{r,\dots,r}_{m+n+1})}(X,Y,z),$
-where $\lambda$ runs over partitions of length at most $n+1$ such that $\lambda_{1}\leq r$, and $(r^{m-n})\cup\lambda$ stands for the partition $(\underbrace{r,\dots,r}_{m-n},\lambda_{1},\dots,\lambda_{n+1})$.
-We write some lemmas before presenting the generalization (see 4.7).
-Recall the Cauchy–Binet formula.
-Given an $m\times n$ matrix $X=(x_{i,j})_{{1\leq i\leq m,1\leq j\leq n}}$ and a subset $B\subset[n]=\{1,\dots,n\}$ of column indices, let $X[B]$ be the submatrix of $X$ obtained by picking up colmns indexed by elements of $B$.
+where $\lambda$ runs over partitions of length at most $n+1$ such that $\lambda_{1}\leq r$, and $(r^{m-n})\cup\lambda$ denotes the partition $(\underbrace{r,\dots,r}_{m-n},\lambda_{1},\dots,\lambda_{n+1})$.
+We recall some lemmas before presenting the generalization (see 4.7).
+We first recall the Cauchy-Binet formula.
+Given an $m\times n$ matrix $X=(x_{i,j})_{{1\leq i\leq m,1\leq j\leq n}}$ and a subset $B\subset[n]=\{1,\dots,n\}$ of column indices, let $X[B]$ be the submatrix of $X$ consisting of the columns indexed by elements of $B$.
 Lemma 4.2.
 For two $m\times n$ matrices $X$ and $Y$, we have
-$\sum_{B}\det X[B]\det Y[B]=\det(XY^{\prime}),$
+$\sum_{B}\det X[B]\det Y[B]=\det(XY^{t}),$
 where the sum runs over $m$-elements subset of $[n]$.
@@
 $\begin{split}[(x_{1}\cdots x_{n})^{r}\mathrm{sp}_{\lambda}(X)]&|_{x_{1}=0}\\ =&\begin{cases}(x_{2}\cdots x_{n})^{r}\mathrm{sp}_{(\lambda_{2},\dots,\lambda_{n})}(x_{2},\dots,x_{n})&\text{ if }\lambda_{1}=r,\\ 0&\text{ otherwise,}\end{cases}\end{split}$
-where $F|_{x_{1}=0}$ means that we substitute $x_{1}=0$ in $F$.
+where $F|_{x_{1}=0}$ is the polynomial obtained by substituting $x_{1}=0$ in $F$.
 We need the following generalization of 4.3 in the proof of 4.7.
@@
 Let $\lambda$ be a partition of length at most $n$ and $\lambda_{1}\leq r$. Then
-$(x_{1}\cdots x_{n})^{r}\operatorname{spo}_{\lambda}(X/z)$
+$(x_{1}\cdots x_{n})^{r}\operatorname{spo}_{\lambda}(X;z)$
 is a polynomial in $x_{1},\dots,x_{n}$, and we have
-$\begin{split}[(x_{1}\cdots x_{n})^{r}\operatorname{spo}_{\lambda}(X/z)]&|_{x_{1}=0}\\ =&\begin{cases}(x_{2}\cdots x_{n})^{r}\operatorname{spo}_{(\lambda_{2},\dots,\lambda_{n})}(x_{2},\dots,x_{n}/z)&\text{ if }\lambda_{1}=r,\\ 0&\text{ otherwise,}\end{cases}\end{split}$
-where $F|_{x_{1}=0}$ means that we substitute $x_{1}=0$ in $F$.
+$\begin{split}[(x_{1}\cdots x_{n})^{r}\operatorname{spo}_{\lambda}(X;z)]&|_{x_{1}=0}\\ =&\begin{cases}(x_{2}\cdots x_{n})^{r}\operatorname{spo}_{(\lambda_{2},\dots,\lambda_{n})}(x_{2},\dots,x_{n};z)&\text{ if }\lambda_{1}=r,\\ 0&\text{ otherwise,}\end{cases}\end{split}$
+where $F|_{x_{1}=0}$ is the polynomial obtained by substituting $x_{1}=0$ in $F$.
 Proof.
 By 1.4, we have
-$\operatorname{spo}_{\lambda}(X/z)=\frac{1}{D}\det\left(x_{i}^{\lambda_{j}+n-j+1}-\bar{x}_{i}^{\lambda_{j}+n-j+1}+z\left(x_{i}^{\lambda_{j}+n-j}-\bar{x}_{i}^{\lambda_{j}+n-j}\right)\right),$
-where $D=\prod_{i=1}^{n}(x_{i}-\bar{x}_{i})\prod_{1\leq i<j\leq n}(x_{i}+\bar{x}_{i}-x_{j}-\bar{x}_{j})$. Let $A_{\lambda}(X)\coloneqq(a_{i,j})_{1\leq i,j\leq n}$ such that
+$\operatorname{spo}_{\lambda}(X;z)=\frac{1}{D}\det\left(x_{i}^{\lambda_{j}+n-j+1}-\bar{x}_{i}^{\lambda_{j}+n-j+1}+z\left(x_{i}^{\lambda_{j}+n-j}-\bar{x}_{i}^{\lambda_{j}+n-j}\right)\right),$
+where $D=\prod_{i=1}^{n}(x_{i}-\bar{x}_{i})\prod_{1\leq i<j\leq n}(x_{i}+\bar{x}_{i}-x_{j}-\bar{x}_{j})$. Let $A_{\lambda}(X)\coloneqq(a_{i,j})_{1\leq i,j\leq n},$ where
 $\begin{split}a_{i,j}=x_{i}^{r+n}\left(x_{i}^{\lambda_{j}+n-j+1}-\bar{x}_{i}^{\lambda_{j}+n-j+1}+z\left(x_{i}^{\lambda_{j}+n-j}-\bar{x}_{i}^{\lambda_{j}+n-j}\right)\right)&\\ =x_{i}^{\lambda_{j}+r+2n-j+1}-x_{i}^{r-\lambda_{j}+j-1}+&z\left(x_{i}^{\lambda_{j}+r+2n-j}-x_{i}^{r-\lambda_{j}+j}\right).\end{split}$
 Then
-$(x_{1}\cdots x_{n})^{r}\operatorname{spo}_{\lambda}(X/z)=\frac{\det A_{\lambda}(X)}{x_{1}^{n}\cdots x_{n}^{n}}\times\frac{1}{D}.$
-Since $\lambda_{j}+r+2n-j+1>\lambda_{j}+r+2n-j>r-\lambda_{j}+j>r-\lambda_{j}+j-1\geq 0,$ $(x_{1}\cdots x_{n})^{r}\operatorname{spo}_{\lambda}(X/z)$ is a polynomial in $x_{1},\dots,x_{n}$.
-Also, note that $r-\lambda_{j}+j-1=0$ if and only if $j=1$ and $\lambda_{1}=r$.
-Therefore,
-$a_{1,j}|_{x_{1}=0}=\begin{cases}-1&\lambda_{1}=r\text{ and }j=1\\ 0&\text{ otherwise.}\end{cases}$
+$(x_{1}\cdots x_{n})^{r}\operatorname{spo}_{\lambda}(X;z)=\frac{\det A_{\lambda}(X)}{x_{1}^{n}\cdots x_{n}^{n}}\times\frac{1}{D}.$
+Since
+$\lambda_{j}+r+2n-j+1>\lambda_{j}+r+2n-j>r-\lambda_{j}+j>r-\lambda_{j}+j-1\geq 0,$
+$(x_{1}\cdots x_{n})^{r}\operatorname{spo}_{\lambda}(X;z)$ is a polynomial in $x_{1},\dots,x_{n}$. Also, note that $r-\lambda_{j}+j-1=0$ if and only if $j=1$ and $\lambda_{1}=r$. Therefore,
+$a_{1,j}|_{x_{1}=0}=\begin{cases}-1&\lambda_{1}=r\text{ and }j=1,\\ 0&\text{ otherwise,}\end{cases}$
 and
@@
 If $\lambda_{1}=r$, then we have
-$\begin{split}[(x_{1}\cdots x_{n})^{r}\operatorname{spo}_{\lambda}(X/z)]|_{x_{1}=0}&=\frac{-(x_{2}\cdots x_{n})\det A_{(\lambda_{2},\dots,\lambda_{n})}(x_{2},\dots,x_{n})}{-(x_{2}\cdots x_{n})^{n}\prod_{i=2}^{n}(x_{i}-\bar{x}_{i})\prod_{2\leq i<j\leq n}(x_{i}+\bar{x}_{i}-x_{j}-\bar{x}_{j})}\\ &=(x_{2}\cdots x_{n})^{r}\operatorname{spo}_{(\lambda_{2},\dots,\lambda_{n})}(x_{2},\dots,x_{n}/z).\end{split}$
+$\begin{split}[(x_{1}\cdots x_{n})^{r}\operatorname{spo}_{\lambda}(X;z)]|_{x_{1}=0}&=\frac{-(x_{2}\cdots x_{n})\det A_{(\lambda_{2},\dots,\lambda_{n})}(x_{2},\dots,x_{n})}{-(x_{2}\cdots x_{n})^{n}\prod_{i=2}^{n}(x_{i}-\bar{x}_{i})\prod_{2\leq i<j\leq n}(x_{i}+\bar{x}_{i}-x_{j}-\bar{x}_{j})}\\ &=(x_{2}\cdots x_{n})^{r}\operatorname{spo}_{(\lambda_{2},\dots,\lambda_{n})}(x_{2},\dots,x_{n};z).\end{split}$
 This completes the proof. ∎
 Lemma 4.5 ([Oka20, Lemma 3.2]).
-Let $p(x,y,z,a,b)$ be a rational function in $x$, $y$, $z$, $a$ and $b$ given by
+Let $p(x,y,z,a,b)$ be the rational function in $x$, $y$, $z$, $a$ and $b$ given by
 $\begin{split}p(x,y,&z,a,b)=\\ &\frac{(1-xz)(1-yz)}{1-xy}-a\frac{(x-z)(1-yz)}{x-y}+b\frac{(1-xz)(y-z)}{x-y}-ab\frac{(x-z)(y-z)}{1-xy}.\end{split}$
@@
 Lemma 4.6.
-Let $q(x,y,z,a,b)$ be a rational function in $x$, $y$, $z$, $a$ and $b$ given by
+Let $q(x,y,z,a,b)$ be the rational function in $x$, $y$, $z$, $a$ and $b$ given by
 (4.2) $\begin{split}q(x,y,&z,a,b)=\\ &\frac{(1+xz)(1+yz)}{1-xy}-a\frac{(x+z)(1+yz)}{x-y}+b\frac{(1+xz)(y+z)}{x-y}-ab\frac{(x+z)(y+z)}{1-xy}.\end{split}$
@@
 Let $m$ and $n$ be positive integers with $n\leq m$, and $r$ a nonnegative integer. Then we have
-(4.4) $\begin{split}\sum_{\lambda}z^{r}\operatorname{spo}_{\lambda}(X/z)\operatorname{spo}_{(r^{m-n})\cup\lambda}&(Y/z)\\ =\sum_{j=1}^{m+n+1}&z^{r+m+n+1-j}\mathrm{sp}_{\big{(}\underbrace{r,\dots,r}_{j-1},\underbrace{r-1,\dots,r-1}_{m+n+1-j}\big{)}}(X,Y),\end{split}$
-where $\lambda$ runs over partitions of length at most $n$ such that $\lambda_{1}\leq r$.
+(4.4) $\begin{split}\sum_{\lambda}z^{r}\operatorname{spo}_{\lambda}(X;z)\operatorname{spo}_{(r^{m-n})\cup\lambda}&(Y;z)\\ =\sum_{j=1}^{m+n+1}&z^{r+m+n+1-j}\mathrm{sp}_{\big{(}\underbrace{r,\dots,r}_{j-1},\underbrace{r-1,\dots,r-1}_{m+n+1-j}\big{)}}(X,Y),\end{split}$
+where $\lambda$ runs over partitions of length at most $n$ such that $\lambda_{1}\leq r$ and $(r^{m-n})\cup\lambda$ denotes the partition $(\underbrace{r,\dots,r}_{m-n},\lambda_{1},\dots,\lambda_{n})$.
 Proof.
@@
 $\begin{cases}X_{i,p}&=\quad x_{i}^{p}-\bar{x}_{i}^{p}+z(x_{i}^{p-1}-\bar{x}_{i}^{p-1}),\\ Y_{i,p}&=\quad y_{i}^{p}-\bar{y}_{i}^{p}+z(y_{i}^{p-1}-\bar{y}_{i}^{p-1}),\end{cases}$
-for $1\leq i\leq n$, $1\leq p\leq M$. Then from a direct but lengthy computation, one can see that the $(i,j)$’th entry of $XY^{\prime}$ is given by
+for $1\leq i\leq n$, $1\leq p\leq M$. By using the formula for the geometric sum, it is immediate that the $(i,j)$-th entry of $XY^{t}$ is given by
 $\bar{x}_{i}^{M}\bar{y}_{j}^{M}q(x_{i},y_{j},z,x_{i}^{2M},y_{j}^{2M}),$
@@
 Therefore,
-(4.5) $\det XY^{\prime}=(\bar{x}_{1}\cdots\bar{x}_{n})^{M}(\bar{y}_{1}\cdots\bar{y}_{n})^{M}\det\left(q(x_{i},y_{j},z,x_{i}^{2M},y_{j}^{2M})\right).$
+(4.5) $\det XY^{t}=(\bar{x}_{1}\cdots\bar{x}_{n})^{M}(\bar{y}_{1}\cdots\bar{y}_{n})^{M}\det\left(q(x_{i},y_{j},z,x_{i}^{2M},y_{j}^{2M})\right).$
 By 4.6, we have
@@
 $\left(W^{(M)}\right)_{i,j}=\begin{cases}x_{i}^{j-1}-x_{i}^{2M+2n+1-j}&1\leq i\leq n\\ y_{i-n}^{j-1}-y_{i-n}^{2M+2n+1-j}&n+1\leq i\leq 2n\\ (-z)^{2n+1-j}&i=2n+1,\end{cases}$
-for $11\leq j\leq 2n+$ 1. Suppose $I_{n}(\lambda)=(\lambda_{1}+n,\dots,\lambda_{n}+1)$. Then by 1.4 and the Cauchy–Binet formula in 4.2, the left hand side of (4.4) is given by
-(4.7) $\begin{split}\sum_{\lambda}z^{r}\operatorname{spo}_{\lambda}(X/z)\operatorname{spo}_{\lambda}(Y/z)=&\displaystyle\sum_{\lambda}\frac{z^{r}}{D_{x}D_{y}}\det X[I_{n}(\lambda)]\,{\det Y[I_{n}(\lambda)]}\\ =\frac{z^{r}}{D_{x}D_{y}}\det(XY^{\prime})=&\frac{(-1)^{n}z^{r}(\bar{x}_{1}\cdots\bar{x}_{n})^{M}(\bar{y}_{1}\cdots\bar{y}_{n})^{M}}{D_{x}D_{y}\displaystyle\prod_{i=1}^{n}\prod_{j=1}^{n}(x_{i}-y_{j})(1-x_{i}y_{j})}\times\det W^{(M)},\end{split}$
+for $11\leq j\leq 2n+$ 1. Suppose $I_{n}(\lambda)={\{}\lambda_{1}+n,\dots,\lambda_{n}+1{\}}$. Then by 1.4 and the Cauchy–Binet formula in 4.2, the left hand side of (4.4) is given by
+(4.7) $\begin{split}\sum_{\lambda}z^{r}\operatorname{spo}_{\lambda}(X;z)\operatorname{spo}_{\lambda}(Y;z)=&\displaystyle\sum_{\lambda}\frac{z^{r}}{D_{x}D_{y}}\det X[I_{n}(\lambda)]\,{\det Y[I_{n}(\lambda)]}\\ =\frac{z^{r}}{D_{x}D_{y}}\det(XY^{t})=&\frac{(-1)^{n}z^{r}(\bar{x}_{1}\cdots\bar{x}_{n})^{M}(\bar{y}_{1}\cdots\bar{y}_{n})^{M}}{D_{x}D_{y}\displaystyle\prod_{i=1}^{n}\prod_{j=1}^{n}(x_{i}-y_{j})(1-x_{i}y_{j})}\times\det W^{(M)},\end{split}$
 where the last equality uses (4.5) and (4.6) and $D_{x}=\prod_{i=1}^{n}(x_{i}-\bar{x}_{i})\prod_{1\leq i<j\leq n}(x_{i}+\bar{x}_{i}-x_{j}-\bar{x}_{j})$, $D_{y}=\prod_{i=1}^{n}(y_{i}-\bar{y}_{i})\prod_{1\leq i<j\leq n}(y_{i}+\bar{y}_{i}-y_{j}-\bar{y}_{j})$.
 Simplifying (4.7), we have
-$\sum_{\lambda}z^{r}\operatorname{spo}_{\lambda}(X/z)\operatorname{spo}_{\lambda}(Y/z)=\frac{(-1)^{r+n}\prod_{i=1}^{n}x_{i}^{n}y_{i}^{n}}{D_{x}D_{y}\displaystyle\prod_{i=1}^{n}\prod_{j=1}^{n}(x_{i}-y_{j})(1-x_{i}y_{j})}\det W_{0},$
+$\sum_{\lambda}z^{r}\operatorname{spo}_{\lambda}(X;z)\operatorname{spo}_{\lambda}(Y;z)=\frac{(-1)^{r+n}\prod_{i=1}^{n}x_{i}^{n}y_{i}^{n}}{D_{x}D_{y}\displaystyle\prod_{i=1}^{n}\prod_{j=1}^{n}(x_{i}-y_{j})(1-x_{i}y_{j})}\det W_{0},$
 where
@@
 Therefore,
-$\begin{split}\sum_{\lambda}z^{r}\operatorname{spo}_{\lambda}(X/z)\operatorname{spo}_{\lambda}(Y/z)=\frac{(-1)^{r+n^{2}+n}}{D_{x}D_{y}\prod_{i=1}^{n}\prod_{j=1}^{n}(x_{i}+\bar{x}_{i}-y_{j}-\bar{y}_{j})}&\det W_{0}\\ =\sum_{j=1}^{2n+1}z^{r+2n+1-j}\mathrm{sp}_{\big{(}\underbrace{r,\dots,r}_{j-1},\underbrace{r-1,\dots,r-1}_{2n+1-j}\big{)}}&(X,Y),\end{split}$
+$\begin{split}\sum_{\lambda}z^{r}\operatorname{spo}_{\lambda}(X;z)\operatorname{spo}_{\lambda}(Y;z)&=\frac{(-1)^{r+n^{2}+n}}{D_{x}D_{y}\prod_{i=1}^{n}\prod_{j=1}^{n}(x_{i}+\bar{x}_{i}-y_{j}-\bar{y}_{j})}\det W_{0}\\ &=\sum_{j=1}^{2n+1}z^{r+2n+1-j}\mathrm{sp}_{\big{(}\underbrace{r,\dots,r}_{j-1},\underbrace{r-1,\dots,r-1}_{2n+1-j}\big{)}}(X,Y),\end{split}$
 where the last equality uses (2.10).
@@
 Now we prove the general case by downward induction on $n$.
-We assume (4.4).
 Multiplying both sides of (4.4) by $(x_{1}\cdots x_{n}y_{1}\cdots y_{m})^{r}$ and then substituting $x_{1}=0$, by 4.3 and 4.4, we have
-$\begin{split}\sum_{\lambda}z^{r}(x_{2}\cdots x_{n})^{r}\operatorname{spo}_{(\lambda_{2},\dots,\lambda_{n})}(x_{2},\dots,x_{n}/z)(y_{1}\cdots y_{m})^{r}\operatorname{spo}_{(r^{m-n})\cup\lambda}(Y&/z)\\ =\sum_{j=2}^{m+n+1}z^{r+m+n+1-j}(x_{2}\cdots x_{n}y_{1}\cdots y_{m})^{r}\mathrm{sp}_{\big{(}\underbrace{r,\dots,r}_{j-2},\underbrace{r-1,\dots,r-1}_{m+n+1-j}\big{)}}&(x_{2},\dots,x_{n},y_{1},\dots,y_{m}),\end{split}$
+$\begin{split}\sum_{\lambda}z^{r}(x_{2}\cdots x_{n})^{r}\operatorname{spo}_{(\lambda_{2},\dots,\lambda_{n})}(x_{2},\dots,x_{n};z)(y_{1}\cdots y_{m})^{r}\operatorname{spo}_{(r^{m-n})\cup\lambda}(Y&;z)\\ =\sum_{j=2}^{m+n+1}z^{r+m+n+1-j}(x_{2}\cdots x_{n}y_{1}\cdots y_{m})^{r}\mathrm{sp}_{\big{(}\underbrace{r,\dots,r}_{j-2},\underbrace{r-1,\dots,r-1}_{m+n+1-j}\big{)}}&(x_{2},\dots,x_{n},y_{1},\dots,y_{m}),\end{split}$
 where the summation on the left hand side is over partitions $\lambda$ of length at most $n$ such that $\lambda_{1}=r$.
 Since $(r^{m-n})\cup\lambda=(r^{m-n+1})\cup(\lambda_{2},\dots,\lambda_{n})$ for such a partition, we have
-$\begin{split}\sum_{\mu}z^{r}\operatorname{spo}_{\mu}(x_{2},\dots,x_{n}/z)\operatorname{spo}_{(r^{m-n+1})\,\cup\,\mu}(Y&/z)\\ =\sum_{j=1}^{m+n}z^{r+m+n-j}\mathrm{sp}_{\big{(}\underbrace{r,\dots,r}_{j-1},\underbrace{r-1,\dots,r-1}_{m+n-j}\big{)}}&(x_{2},\dots,x_{n},y_{1},\dots,y_{m}),\end{split}$
+$\begin{split}\sum_{\mu}z^{r}\operatorname{spo}_{\mu}(x_{2},\dots,x_{n};z)&\operatorname{spo}_{(r^{m-n+1})\,\cup\,\mu}(Y;z)\\ =\sum_{j=1}^{m+n}&z^{r+m+n-j}\mathrm{sp}_{\big{(}\underbrace{r,\dots,r}_{j-1},\underbrace{r-1,\dots,r-1}_{m+n-j}\big{)}}(x_{2},\dots,x_{n},y_{1},\dots,y_{m}),\end{split}$
 where $\mu$ runs over all partitions of length at most $n-1$ such that $\mu_{1}\leq r$. This completes the proof. ∎
+Declarations
+The authors have no competing interests to declare that are relevant to the content of this article.
 Acknowledgements
-We thank A. Ayyer and A. Stokke for very helpful discussions.
+We thank A. Ayyer and A. Stokke for very helpful discussions. We acknowledge support from the Department of Science and Technology grant CRG/2021/001592.